### v1

# MPP-2025-122

Abstract
To give more credence to the M-theoretic Emergence Proposal it is important to show that also classical kinetic terms in a low energy effective action arise as a quantum effect from integrating out light towers of states. We show that for compactifications of type IIA on Calabi-Yau manifolds, the classical weak coupling Yukawa couplings, which are the triple intersection numbers of the Calabi-Yau threefold, can be obtained from the 1/2-BPS protected one-loop Schwinger integral over $D2$-$D0$ bound states, after employing a novel regularization for the final infinite sum of Gopakumar-Vafa invariants. Approaching the problem in a consecutive manner from 6D decompactification over emergent string to the ultimate M-theory limits, we arrive at a mathematically concrete regularization that involves finite distance degeneration limits of Calabi-Yau threefolds in an intriguing way. We test and challenge this proposal by the concrete determination of the periods around such degeneration points for threefolds with one Kähler modulus and the two examples $\mathbb{P}_{1,1,1,6,9}[18]$ and $\mathbb{P}_{1,1,2,2,6}[12]$.
1 Introduction
The Emergence Proposal [1, 2, 3] is among the most promising and yet less understood ideas of the swampland program, aiming at characterizing universal features of gravitational effective theories. Its strongest version [4] holds that the dynamics (kinetic terms) for all fields are emergent in the infrared by integrating out towers of states down from an ultraviolet scale, which is below the Planck scale. The species scale [5, 6] provides a natural candidate for such an ultraviolet scale, which has been employed in the subsequent field theory approaches of [7, 8, 9, 10].
However, as long as one sticks to this strong version of the Emergence Proposal, there seems to be little room to implement it exactly in full string theory, see e.g. [11, 12, 13, 14]. In the spirit of the Emergent String Conjecture [15], the only other candidate are decompactification limits. And indeed, given our ignorance of M-theory and the quantum origin of graviton scattering in the BFSS matrix model, it is a logical possibility that the (strong) Emergence Proposal might be realized therein. This was first suggested in [14] and then elaborated on in a series of papers [16, 17, 18, 19]. In particular, in [17] the precise form of the $R^{4}$ interaction in M-theory, including the correct relative coefficients, was recovered by integrating out at one-loop level infinite towers of states with typical mass not larger than the eleven dimensional Planck scale, which is the species scale in the studied limit and duality frame. This computation was generalized to a higher derivative $F^{4}$ interaction in type IIA compactified on a $K3$ to six dimensions [19].
However, as it is clear from its initial formulation, the Emergence Proposal is first and perhaps foremost about kinetic terms, so one should be able to recover them by integrating out towers of states. This is challenging in general, especially in a gravitational context. For example, due to the equivalence principle, the Einstein-Hilbert term shall in principle receive contributions from all states in the theory, for gravity couples to everything. To simplify the task, one can look at less realistic kinetic interactions, possibly with some protection under a certain symmetry. A good example of such an interaction can be found in Calabi-Yau (CY) compactifications of the type IIA superstring leading to a four-dimensional theory with eight preserved supercharges. Here, kinetic terms are encoded into a holomorphic function of the Kähler moduli, the prepotential. More specifically, the tree-level contribution at weak type IIA string coupling is encoded in a cubic polynomial with coefficients related to the triple intersection numbers (TINs) of the CY. Hence, if the M-theoretic Emergence Proposal is to be correct, then one should be able to reproduce all TINs by integrating out the light infinite towers of states in a co-scaled strong coupling limit of type IIA, where one scales all Kähler moduli isotropically such that the 4D Planck scale remains fixed. Henceforth, we will refer to this limit as the (isotropic) M-theory limit. As generally shown in [16], the light towers with a mass scale below the species scale, i.e. the eleven dimensional Planck-scale, are transverse $M2$- and $M5$-branes carrying KK momentum along the eleventh (type IIA $D0$-branes) and any other compact dimension.
A first step in this direction was performed in [14] for a non-compact CY, namely the resolved conifold, where the geometry only admits a single 2-cycle, on which an $M2$-brane could be wrapped. After realizing that the M-theory amplitude is nothing else than the Schwinger integral of Gopakumar/Vafa [20, 21], which is precisely integrating out 1/2 BPS bound states of $D2$- and $D0$-branes, a proper regularization of its UV-divergence followed by $\zeta$-function regularization also gave the correct formal TIN of the resolved conifold. It is the purpose of this paper to generalize this emergence computation of TINs to compact CY threefolds, where finally one has to sum over the full homology lattice $H_{2}(X,\mathbb{Z})$ weighted by the exponentially growing Gopakumar-Vafa (GV) invariants [^1] Let us notice that an alternative approach has been put forward in a series of recent papers by Hattab/Palti [22, 23, 24]. In particular, the initial real Schwinger integral of Gopakumar/Vafa has been changed to a complex contour integral.. It seems to be a horrendous task to systematically define a regularization of these highly divergent infinite sums that is guaranteed to give the correct TINs.
For appreciating the approach we take, it is important to recall that in the former emergence computations [14, 16, 17, 18, 19], the one-loop amplitude was initially defined via an integral over a real Schwinger parameter, similar to naive string amplitudes before implementing level matching and modular invariance. Such integrals are notoriously UV-divergent, but in [14, 17] a way to regularize them was proposed, which involved a minimal substraction of the divergent terms followed by a $\zeta$-function regularization of the final infinite sums. It was shown in [17] that this procedure also gave the correct finite results upon applying it to one-loop string amplitudes, which constitutes a non-trivial consistency check. The question is whether a generalization of this regularization can be found that produces finite TINs for infinite sums of the type
$\begin{split}Y^{(0)}_{t_{i}t_{j}t_{k}}={1\over 2}\sum_{\beta\in H_{2}(X,\mathbb{Z})}\alpha_{0}^{\beta}\,\beta_{i}\beta_{j}\beta_{k}\,,\end{split}$ (1.1)
where $\alpha_{0}^{\beta}$ denote the Gopakumar-Vafa invariants which are known to generically grow as $\alpha_{0}^{\beta}\sim\exp(\gamma\beta)$ with $\gamma>0$ [25].
This paper is organized as follows: In section 2 we explain in more detail the origin of these infinite sums and provide a mathematical toy example, which already contains an essential hint of how one could approach this problem. Before directly dealing with the most involved isotropic M-theory limits, in section 3 we first consider a couple of special CYs, which either admit an elliptic or a $K3$-fibration. For the elliptic fibrations we take a 6D decompactification limit so that only $M2$-branes wrapping the elliptic fiber are among the light towers of states. Integrating them out can be performed just via $\zeta$-function regularization and precisely yields the correct constant to complete the modular form $E_{4}(t)$, where $t$ measures the size of the fiber. This serves as a first hint that indeed the regularization of infinite sums of the type (1.1) leads to sensible results. For the $K3$-fibration we take the emergent string limit which admits a dual weakly coupled heterotic string description. In this case one expects that only the triple intersection numbers not involving the base 2-cycle are emerging from integrating out $D2$-$D0$ bound states. Employing the heterotic dual we can determine the relevant GV invariants and concretely formulate and test a proposal of how the regularization of (1.1) can be performed. This involves a $t\to 0$ limit and modular S-transformation in an intriguing way.
Building on these working examples, in section 4 we approach the actual problem of the emergence of the complete set of TINs in isotropic M-theory limits. As will become clear this part is more speculative, as the aforementioned limit $t\to 0$ cannot be taken. This is first discussed for the Quintic threefold, where the prepotential already diverges at the location of the conifold singularity, which is correlated with the exponential growth of the GV invariants. The analysis of this example will guide us to a mathematically concrete regularization of the infinite sum (1.1) that involves the wanted TIN and the limit of the known prepotential in the type IIA weak coupling regime towards degeneration limits of the CY. For the Quintic and the other 13 CY threefolds from [26] with $h_{11}=1$ this gives the correct TIN from the limit to the location of the conifold singularity.
In the rest of the paper we challenge this admittedly speculative proposal for two CYs having two Kähler moduli, namely the elliptic fibration $\mathbb{P}_{1,1,1,6,9}[18]$ featuring two intersecting conifold loci and the $K3$-fibration $\mathbb{P}_{1,1,2,2,6}[12]$ having two intersection points of a conifold and another degeneration locus, often called (heterotic dual) strong coupling singularity. We will find that for conifold loci our proposal works quite straightforwardly while the other degeneration locus reveals some subleties that require a further refinement of the regularization procedure.
On the more technical side, these computations involve the knowledge of the CY periods in the vicinity of these degeneration loci in the complex structure moduli space [^2] We are indebted to Rafael Álvarez-García for sharing his insights into these computations and for actually providing some of the results and teaching us how to achieve them., i.e. results that are not generally available in the literature, yet. The details of these computations will be reported in [27]. We also add two appendices. In the first, we review the relation between the asymptotic growth of GV invariants and the appearance of degeneration loci in the complex structure moduli space and also mention a few observations once one tries to generalize the original computation for the Quintic [28] to the CY $\mathbb{P}_{1,1,2,2,6}[12]$. In the second appendix we point out that in case of a description of the CY in terms of a Gauged Linear Sigma Model (GLSM), the location of the singularity is nothing else than the quantum corrected singular point in the middle of the classical phase diagram of this GLSM.
2 Preliminaries
In [14] it has been suggested that the strong Emergence Proposal might be realized in M-theory by looking at compactifications of type IIA string theory on a CY threefold $X$ yielding $N=2$ supergravity in four dimensions. The corresponding vector-multiplet moduli space is spanned by real scalars $\tau_{i}$ which, together with the Kalb-Ramond axions $b_{i}$, define the complex Kähler moduli $t_{i}=b_{i}+i\tau_{i}$. The kinetic terms and gauge couplings of the vector-multiplets are determined by a holomorphic prepotential $\mathcal{F}_{0}(t)$. Due to supersymmetry, this is tree-level exact in type IIA and reads [^3] In this work we will omit writing the linear and quadratic terms, whose prefactors can be altered by acting with symplectic transformations [29]. However, part of that information is important for fixing an integral symplectic basis for the periods [30, 31].
$\mathcal{F}_{0}(t)=\frac{(2\pi i)^{3}}{g_{s}^{2}}\Big{[}\frac{1}{6}\kappa^{ijk}\,t_{i}t_{j}t_{k}-\frac{\zeta(3)}{2(2\pi i)^{3}}\chi(X)+{1\over(2\pi i)^{3}}\sum_{\beta\in H_{2}(X,\mathbb{Z})}\alpha_{0}^{\beta}\,{\rm Li}_{3}(e^{2\pi i\beta\cdot t})\Big{]}\,,$ (2.1)
where $\kappa^{ijk}=\kappa_{t_{i}t_{j}t_{k}}$ are the TINs of $X$, $\chi(X)$ its Euler characteristic and $\alpha_{0}^{\beta}$ are genus zero GV invariants [20, 21].
It is the purpose of this paper to investigate whether one can obtain the classical cubic term in the prepotential from a one-loop Schwinger-like integral in the strong coupling regime, i.e. the isotropic M-theory limit. Following [14], it is defined by scaling all type IIA Kähler moduli as $\tau_{i}\to\lambda\,\tau_{i}$ and $g_{s}\to\lambda^{3/2}g_{s}$ (with $\lambda\to\infty$) so that the 4D Planck scale remains constant. Then, the $D0$-branes are the lightest species leading to the species scale $\Lambda_{\rm sp}\sim M_{\rm pl}/\lambda^{1/2}$, which is at threshold with the mass of $D2$-branes wrapping 2-cycles of the CY. This is a decompactification limit to 5D described by M-theory on the CY. When working in M-theory units, this co-scaled limit leaves constant all radii of the CY measured in units of the 11D Planck scale $M_{*}$ and scales $r_{11}\to\lambda\,r_{11}$, $M_{*}\to M_{*}/\lambda^{1/2}$. Then, the species scale is nothing else than the 5D Planck scale, scaling in the same way as $M_{*}$.
Integrating out the light towers of states with mass scale below or at the species scale we arrive at the prescription of Gopakumar/Vafa [20, 21], i.e. we need to integrate out wrapped $M2$-branes transverse with respect to the M-theory circle but with KK-momentum along it. In type IIA variables, these are $D2$-$D0$ bound states. The corresponding Schwinger integral reads
$\mathcal{F}_{0}(t)=\sum_{(\beta,n)\neq(0,0)}\alpha_{0}^{\beta}\int_{0}^{\infty}\frac{ds}{s^{3}}\;e^{sZ_{n}(\beta)}\,,$ (2.2)
where $Z_{n}(\beta)=\frac{2\pi i}{g_{s}}(\beta\cdot t-n)$ is the central charge of the supersymmetry algebra. The complete (perturbative) $\mathcal{F}_{0}$ for the (non-compact) resolved conifold, which has only a single shrinkable 2-cycle and therefore no sum over $\beta$, has been successfully reproduced in [14]. There, the divergent sum over the number of $D0$-branes was regularized with the analytic continuation of the Riemann $\zeta$-function. For a general compact CY, the $D2$-$D0$ bound states wrap all 2-cycles in $H_{2}(X,\mathbb{Z})$ and therefore another infinite sum needs to be performed.
2.1 Quest for emergence of kinetic terms
To set the stage, let us focus for the moment on compact threefolds with $h_{11}=1$ and a single TIN $\kappa$. We can extract information about the cubic term from (2.2) by taking the third derivative with respect to $t$, giving the “Yukawa coupling”
[…]
After some steps, we can rewrite it as
$Y_{ttt}=\sum_{\beta\neq 0}\alpha_{0}^{\beta}\,\beta^{3}\left(\frac{1}{2}+\frac{e^{2\pi i\beta t}}{1-e^{2\pi i\beta t}}\right)\,,$ (2.4)
where the second term in the brackets is related to the non-perturbative terms in (2.1), i.e. the ${\rm Li}_{3}$-terms. Note that the factor $1/2$ is precisely the formal triple intersection number for the single two-cycle of the resolved conifold. The whole expression, including the factor 1/2, is reminiscent of the thermal energy-density of a gas of particles in four dimensions with energy levels $\beta$, degeneracy of states $\alpha_{0}^{\beta}\beta^{2}$ and temperature $T=i/t$. The first term in the brackets is then the (divergent) zero point energy, which in the following we call the zero point Yukawa coupling denoted as $Y_{ttt}^{(0)}$. This object is the focus of our interest.
Typically the (genus-zero) GV invariants scale exponentially for large $\beta$, $\alpha_{0}^{\beta}\sim\exp(\gamma\beta)$, with $\gamma>0$. However, for $h_{11}>1$ there can also be sub-towers of wrapped $D2$-branes leading to a weaker scaling such as $\alpha_{0}^{\beta}\sim\exp(\gamma\sqrt{\beta})$, or even to a polynomial behaviour as for KK towers. One can then take a corresponding infinite distance limit so that only these towers contribute. The degeneracy of $\alpha_{0}^{\beta}$ is then closely tied to the type of limit as shown in table 1, where we listed the limits relevant for the following discussion.
limit degeneracy of $\alpha_{0}^{\beta}$
M-theory $\exp(\gamma\beta)$
emergent string $\exp(\gamma\sqrt{\beta})$
6D decompactification $1$
Table 1: Infinite distance limits and degeneracy of genus-zero GV invariants.
In any case, the first sum in (2.4) will be divergent regardless of these limits. Hence, emergence means that upon regularization, the $t$-independent first term in the bracket in (2.4) should give rise to the self-intersection number
$Y_{ttt}^{(0)}:=\frac{1}{2}\sum_{\beta=1}^{\infty}\alpha_{0}^{\beta}\,\beta^{3}\Big{|}_{\rm reg}\stackrel{{\scriptstyle!}}{{=}}\kappa_{ttt}\,.$ (2.5)
Note that here the left-hand side is a quantity to be computed in the strong coupling regime, $g_{s}\gg 1$, i.e. within M-theory, whereas the right hand side is the usual integer-valued tree-level Yukawa coupling arising at weak string coupling, $g_{s}\ll 1$.
Therefore, the quest is to define a regularization such that the zero-point Yukawa coupling, $Y^{(0)}_{ttt}$, matches with the classical contribution in the perturbative string regime. Recall that $Y^{(0)}_{ttt}$ diverges due to the sum over wrapped $D2$-$D0$ bound states in the M-theory regime where, in contrast to string perturbation theory, we currently have no formalism to perform the integral and get to a finite result right away. Indeed, such a computation shall be part of the microscopic description of M-theory, which is lacking at present. Therefore, here we can only head for a properly working regularization scheme, like that of minimal subtraction and $\zeta$-function regularization proposed for the computation of the $R^{4}$ terms in M-theory in [17].
For towers of $D2$-branes with a polynomial scaling of the GV invariants, we can still apply $\zeta$-function regularization, but for the exponentially degenerate towers other methods need to be developed. Our approach is to move from simpler to more complicated cases and in the process refine and adapt the regularization method. Let us start by presenting a regularization method for infinite sums that already demonstrates our main strategy.
2.2 Regularization via modular forms
Suppose we are dealing with sums of the form
$S_{n}=\sum_{k=1}^{\infty}k^{n}\,,$ (2.6)
which are usually regularized with the $\zeta$-function $\zeta(-n)$. It is instructive to recall how this works in practice. One introduces a regulator $q=\exp(-2\pi/\Lambda)$, with $\Lambda\gg 1$, such that for finite $\Lambda$ the sum is finite. Then, one takes the limit $\Lambda\to\infty$ to isolate the divergence. In practice, the conventional way is to introduce $q$ such that derivatives of the geometric series arise, namely
$S_{n}=\lim_{\Lambda\rightarrow\infty}\sum_{k=1}^{\infty}k^{n}q^{k}=\lim_{\Lambda\rightarrow\infty}\left[(q\,\partial_{q})^{n}\left(\frac{q}{1-q}\right)\right]\,,$ (2.7)
for $|q|<1$. One can then expand the right-hand side for large $\Lambda$ to obtain
$S_{n}=\lim_{\Lambda\rightarrow\infty}\left[n!\frac{\Lambda^{n+1}}{(2\pi)^{n+1}}+C_{n}+\mathcal{O}(\Lambda^{-1})\right].$ (2.8)
The divergence appears only at $(n+1)$-th order and can be minimally subtracted, leaving us with the regularized value $C_{n}\equiv\zeta(-n)$.
Let us now arrive at the same result with an alternative method that exploits properties of modular forms to extract the singular term. We assume $n$ to be odd in what follows. The idea is to express the above sum in terms of an Eisenstein series, which requires us to introduce the regulator $q$ in a different fashion, namely
$S_{2n-1}=-2\lim_{\Lambda\rightarrow\infty}\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}k^{2n-1}\,q^{lk}\,.$ (2.9)
The normalization $(-2)$ is chosen in order to compensate with the factor $\zeta(0)=-1/2$ arising from the sum over $l$ in the limit $q\rightarrow 1$. We define the summation index $m=lk$ and then replace the second sum with a sum over the divisors of $m$, giving
$S_{2n-1}=-2\lim_{\Lambda\rightarrow\infty}\sum_{m=1}^{\infty}\sigma_{2n-1}(m)\,q^{m}\qquad\text{with}\quad\sigma_{n}(m)=\sum_{k|m}k^{n}\,.$ (2.10)
This form of $S_{2n-1}$ can now be directly related to the Eisenstein series
$E_{2n}(\tau)=1+c_{2n}\sum_{m=1}^{\infty}\sigma_{2n-1}(m)\,q^{m}\,,\quad$ (2.11)
with
$c_{2n}=\frac{(2\pi i)^{2n}}{(2n-1)!\,\zeta(2n)}=\frac{2}{\zeta(1-2n)}\,.$ (2.12)
In this way, we arrive at
$S_{2n-1}=\lim_{\Lambda\rightarrow\infty}\left[-\frac{2}{c_{2n}}\left(E_{2n}\left({\frac{i}{\Lambda}}\right)-1\right)\right]\,.$ (2.13)
These Eisenstein series are modular forms of degree $2n$, so that under a modular $S$-transformation they transforms as
$E_{2n}\left(-\frac{1}{\tau}\right)=\tau^{2n}E_{2n}(\tau)\,.$ (2.14)
By applying this transformation rule to the first term in (2.13), we find the behaviour
$E_{2n}\left(\frac{i}{\Lambda}\right)=(-1)^{n}\Lambda^{2n}+\mathcal{O}(e^{-2\pi\Lambda})\,.$ (2.15)
A couple of comments are in order here. First, the singularity that has just been isolated has the same degree as the one occurring in the former example with a slightly different regularization scheme. Second, after minimally subtracting this divergence, the leading term is already exponentially suppressed in $\Lambda$ and from the Eisenstein series one does not get any constant contribution surviving the limit $\Lambda\to\infty$. Hence, the only such contribution comes from the second term in (2.13) and gives the regularized value
$S_{2n-1}=\frac{2}{c_{2n}}=\zeta(1-2n)\,.$ (2.16)
For our purposes, the advantage of this regularization scheme is that it allows to explicitly isolate the divergence from the constant term and from the contributions vanishing in the limit $\Lambda\to\infty$. We will see that this method, and a generalization thereof, can be successfully applied to regularize the sum over GV invariants as in (2.5).
3 Large base limits of fibered CY threefolds
Before we discuss the isotropic M-theory limit, where the full homology lattice contributes to the Schwinger integral, we look at simpler cases where only part of it contributes. This restriction is of course only allowed if it is correlated with taking an appropriate infinite distance limit, in which the left out GV invariants give rise to heavy towers of $D2$-branes with mass scale above the species scale. Even though for such limits one will not have one-loop emergence of the full prepotential, integrating out the light towers of $D2$-branes will provide part of it and one can still approach the problem of what the meaning of the zero point Yukawa coupling is.
3.1 Elliptically fibered CY
First, we consider CY threefolds with a sublattice of GV invariants that essentially grow like KK towers (see e.g. the examples in [32]). For such cases, we want to identify an infinite distance limit such that all wrapped $D2$-brane states with exponential degeneracy are among the heavy states that, due to our philosophy of perturbative QG theories, are not integrated out. As we will see, this allows us to carry out the sum over the simple homology lattice by just using $\zeta$-function regularization.
The CY manifolds that feature the desired property are elliptically fibered. As the maybe simplest example, we consider the CY threefold obtained by resolving the $\mathbb{Z}_{3}$ singularity of a degree 18 hypersurface in $\mathbb{P}_{1,1,1,6,9}[18]$ with Hodge numbers $(h_{21},h_{11})=(272,2)$. The GLSM data specifying the resolved threefold are given in table 2 and can be found in the appendix of [30].
$z_{1}$ $z_{2}$ $z_{3}$ $z_{4}$ $z_{5}$ $z_{6}$ $p$
$l^{(1)}$ 0 0 0 2 3 1 -6
$l^{(2)}$ 1 1 1 0 0 -3 0
Table 2: Data specifying elliptically fibered CY
[…]
Defining $H=3L+E$, the intersection numbers of $(H,L)$ were discussed in [33] and read
$\begin{split}H^{3}=9\,,\quad H^{2}\cdot L=3\,,\quad H\cdot L^{2}=1\,,\quad L^{3}=0\,.\end{split}$ (3.1)
The Poincaré dual two-forms are denoted by $(w_{H},w_{L})$ so that we expand the Kähler form as $J=\tau_{1}w_{H}+\tau_{2}w_{L}$. Hence, for the total volume we obtain
$\begin{split}\mathcal{V}_{6}\sim\int J^{3}=9\tau_{1}^{3}+9\tau_{1}^{2}\tau_{2}+3\tau_{1}\tau_{2}^{2}\,.\end{split}$ (3.2)
Moreover, the fiber class $f\in H_{2}(X,\mathbb{Z})$ is given by $L\cap L$ and one can easily show that
$\begin{split}{\cal V}_{2}(f)=\int_{L\cap L}J=\tau_{1}\,.\end{split}$ (3.3)
The second Mori-cone generator is given by the curve $l=L\cap E$ of volume
$\begin{split}{\cal V}_{2}(l)=\int_{l=L\cap(H-3L)}J=\tau_{2}\,.\end{split}$ (3.4)
3.1.1 Integrating out lightest towers of states
Our goal is to take an asymptotic limit where only those $D2$-branes wrapping the elliptic fiber $f$ are considered as perturbative states, since only these GV invariants are constant in this particular case, namely
$\begin{split}\alpha_{0}^{(n_{1},0)}=540\qquad{\rm for}\quad n_{1}\geq 1\,,\end{split}$ (3.5)
which is equal to minus the Euler characteristic of this CY. This can be achieved by the non-isotropic rescaling $\tau_{2}\rightarrow\lambda\tau_{2}$, $g_{s}\to\lambda g_{s}$ with $\lambda\rightarrow\infty$ and $\tau_{1}$ fixed, so that $M_{{\rm pl},4}$ stays fixed. Hence a $D2$-brane wrapping an effective curve $C=n_{1}f+n_{2}l$ bound to $D0$-branes is 1/2 BPS and has mass
$\begin{split}m_{D0-D2}=\frac{M_{s}}{g_{s}}\big{|}n_{1}t_{1}+n_{2}t_{2}+m\big{|}=\frac{M_{{\rm pl},4}}{\sqrt{V_{6}}}\big{|}n_{1}t_{1}+n_{2}t_{2}+m\big{|}\,\end{split}$ (3.6)
with $t_{i}=b_{i}+i\tau_{i}$. Since $V_{6}\sim\lambda^{2}$, it is obvious that the lightest towers of states are $D0$-branes and $D2$-branes wrapping solely the class $f$, i.e. having $n_{2}=0$ and mass
$\begin{split}m\sim\frac{M_{{\rm pl},4}}{\lambda}\,.\end{split}$ (3.7)
Since $\alpha_{0}^{(n_{1},0)}=540$ is constant, we are dealing with two light multiplicative KK towers [34], which yield the species scale
$\begin{split}\Lambda_{\rm sp}=M_{{\rm pl},4}^{\frac{1}{2}}\Delta m^{\frac{1}{2}}\sim\frac{M_{{\rm pl},4}}{\lambda^{1/2}}\,.\end{split}$ (3.8)
In the classification of [15] this limit translates into a $J$-Class A limit in the vector moduli space of M-theory eventually leading to a further decompactification to six dimensions (this limit was previously called type IIIc in [35] and also studied in [36]). However, among the perturbative states there is also a light string, given by an $N\!S5$-brane wrapping the divisor $L$. One can show that there is no 4-cycle with a volume scaling like $\tau_{1}^{2}\sim\lambda^{0}$ and the divisor volumes scale like
$\begin{split}{\cal V}_{4}(H)=\int_{H}J\wedge J=(3\tau_{1}+\tau_{2})^{2}\sim\lambda^{2}\,,\quad{\cal V}_{4}(L)=\int_{L}J\wedge J=\tau_{1}(3\tau_{1}+\tau_{2})\sim\lambda\,.\end{split}$ (3.9)
For the tension of the $N\!S5$ wrapped on $L$ one finds $T_{NS5}=\frac{M_{s}^{2}}{g_{s}^{2}}{\cal V}_{4}(L)$ and hence $M_{\rm str}=\frac{M_{s}}{g_{s}}\sqrt{{\cal V}_{4}(L)}\sim\frac{M_{{\rm pl},4}}{\lambda^{1/2}}$. Therefore, we are not in the isotropic M-theory limit but in a limit where the light towers are string excitations and parametrically lighter bound states of $D0$- and $D2$-branes (wrapped on $f$) signalling a decompactification to 6D.
Using formula (2.4), we can now explicitly integrate out the light $D2$-$D0$ bound states which gives
$\begin{split}Y_{t_{1}t_{1}t_{1}}&=\sum_{k=1}^{\infty}\alpha_{0}^{(k,0)}\,k^{3}\left({1\over 2}+{q^{k}\over(1-q^{k})}\right)=\frac{540}{2}\sum_{k=1}^{\infty}k^{3}+540\sum_{k=1}^{\infty}{k^{3}\,q^{k}\over(1-q^{k})}\\[2.84544pt] &={9\over 4}\left(1+240\sum_{k=1}^{\infty}{k^{3}\,q^{k}\over(1-q^{k})}\right)={9\over 4}E_{4}(t_{1})\,,\end{split}$ (3.10)
with $q=\exp(2\pi it_{1})$ and where we used $\sum_{k=1}^{\infty}k^{3}=\zeta(-3)=1/120$. From this calculation we do not obtain the triple-intersection number $\kappa_{t_{1}t_{1}t_{1}}=9$, but since we are not taking the M-theory limit with isotropic co-scaling of all compact directions, it was also not expected a priori. However, we see that the divergent zero-point Yukawa coupling, after $\zeta$-function regularization provides precisely the constant term in $E_{4}(t_{1})$ that makes it a modular form of degree 4 (under modular transformation of $t_{1}$, the size of the toroidal fiber). While the appearance of a modular form is expected for such an elliptic fibration, the actual point here is to show how this result arises in the non-isotropic large coupling limit from integrating out the light towers of states. This we take as a first indication that also for compact CYs the regularized zero-point Yukawa coupling yields physically meaningful results. Note that for getting this result, the generic factor $1/2$ in the first line of (3.10) was essential.
This computation is expected to generalize to every torus-fibered CY. One example is the CY $\mathbb{P}_{1,1,1,3,6}[12]$ with $(h_{21},h_{11})=(165,3)$, which is a fibration over $\mathbb{P}^{2}$ with two sections. The GV invariants that count $D2$’s only wrapping the fiber are given by $\alpha_{0}^{(2k,0,0)}=324=-\chi(X)$ and $\alpha_{0}^{(2k-1,0,0)}=216$ with $k\geq 1$. For the relevant Yukawa coupling we obtain
$\begin{split}Y_{t_{1}t_{1}t_{1}}&=\sum_{k\geq 1}k^{3}\alpha_{0}^{(k,0,0)}\left({1\over 2}+{q^{k}\over(1-q^{k})}\right)\\ &=216\sum_{k\geq 1}k^{3}\left({1\over 2}+{q^{k}\over(1-q^{k})}\right)+4\cdot 216\sum_{k\geq 1}k^{3}\left({1\over 2}+{q^{2k}\over(1-q^{2k})}\right)\\ &={9\over 10}\Big{(}E_{4}(t_{1})+4E_{4}(2t_{1})\Big{)}\,,\end{split}$ (3.11)
which is a modular form of weight four for $\Gamma^{+}_{0}(2)$. The latter is the extension of the congruence subgroup $\Gamma_{0}(2)$ by the Fricke involution $\omega_{2}$, where $\omega_{N}$ (with $N\in\mathbb{N}$) is defined as
$\begin{split}\omega_{N}=\left(\begin{matrix}0&-{1\over\sqrt{N}}\\ \sqrt{N}&0\end{matrix}\right)\in PSL(2,\mathbb{R})\end{split}$ (3.12)
and induces the transformation $\tau\to-1/(N\tau)$ (see e.g. [37]). For torus fibrations with more sections, the cyclic properties of the GV invariants become more involved. For instance, the CY $\mathbb{P}_{1,1,1,3,3}[9]$ with $(h_{21},h_{11})=(112,4)$ has three sections and the GV invariants with non-trivial fiber class are given by $\alpha_{0}^{(3k,0,0)}=216$ and $\alpha_{0}^{(3k-1,0,0)}=\alpha_{0}^{(3k-2,0,0)}=162$ with $k\geq 1$. Nevertheless, the computation works similarly and yields
$\begin{split}Y_{t_{1}t_{1}t_{1}}=\frac{27}{40}\Big{(}E_{4}(t_{1})+9E_{4}(3t_{1})\Big{)}\,,\end{split}$ (3.13)
which is now a modular form of weight four for $\Gamma^{+}_{0}(3)$.
3.2 $K3$-fibered CY
Our next example is a $K3$-fibration over the base $\mathbb{P}^{1}$. More specifically, we study the CY threefold obtained from a hypersurface of degree 24 in $\mathbb{P}_{1,1,2,8,12}[24]$ with the singularities resolved. The model is known to be dual to the heterotic string on $K3\times T^{2}$ in the limit where the base $\mathbb{P}^{1}$ grows large [38]. We will use this dictionary to extract explicit expressions for the Gopakumar-Vafa invariants of 2-cycles in the $K3$-fiber. Concerning our conventions, we closely follow [39] (see also [40]).
In table (3) we provide the GLSM data for the resolution $\mathbb{P}_{1,1,2,8,12}[24]$ from [30].
$z_{1}$ $z_{2}$ $z_{3}$ $z_{4}$ $z_{5}$ $z_{6}$ $z_{7}$ $p$
$l^{(1)}$ 3 2 0 0 0 1 0 $-6$
$l^{(2)}$ 0 0 0 1 1 0 $-2$ 0
$l^{(3)}$ 0 0 1 0 0 $-2$ 1 0
Table 3: Data specifying $K3$-fibered CY
We denote the type IIA complexified Kähler moduli by $t_{i}$ $(i=1,2,3)$, where $t_{2}$ measures the size of the base. In the limit of large base $t_{2}\rightarrow\infty$ the prepotential reads [^4] Here we have set the prefactor $(2\pi i)^{3}/g_{s}^{2}=1$.
$\begin{split}\mathcal{F}_{0}^{\rm{IIA}}=\mathcal{F}_{0}^{\rm IIA,cubic}+\frac{1}{(2\pi i)^{3}}\sum_{n_{1},n_{3}=0}^{\infty}\alpha_{0}^{(n_{1},0,n_{3})}\,{\rm Li}_{3}(q_{1}^{n_{1}}q_{3}^{n_{3}})+\mathcal{O}\left(e^{2\pi it_{2}}\right)\,,\end{split}$ (3.14)
where world-sheet instantons wrapping also the base become negligible. The cubic term reads
$\begin{split}\mathcal{F}_{0}^{\rm IIA,cubic}=\frac{4}{3}t_{1}^{3}+t_{1}^{2}t_{2}+2t_{1}^{2}t_{3}+t_{1}t_{2}t_{3}+t_{1}t_{3}^{2}\end{split}$ (3.15)
and provides the classical (tree-level) contribution in the weakly coupled type IIA theory with $g_{s}^{\rm IIA}\ll 1$.
This type IIA model is dual to the heterotic $STU$-model, which has been studied in [41]. Here $S$ denotes the (one-loop corrected) axio-dilaton and $(T,U)$ are the Kähler and complex structure moduli of the $T^{2}$. These three heterotic moduli are related to complexified Kähler moduli on the type IIA side as
$\begin{split}t_{1}=iU\,,\quad t_{2}=iS\,,\quad t_{3}=i(T-U)\,.\end{split}$ (3.16)
The Kähler cone is ${\rm Im}(t_{i})>0$ which means that on the heterotic side we are in the chamber ${\rm Re}(T)>{\rm Re}(U)$. The prepotential is given by
$\begin{split}\mathcal{F}_{0}^{\rm het}=i\bigg{(}-\mathcal{F}_{0}^{\rm het,cubic}-\frac{1}{(2\pi)^{3}}&\sum_{k,l=0}^{\infty}2c_{1}(kl){\rm Li}_{3}(e^{-2\pi(kT+lU)})\\[2.84544pt] &-\frac{2}{(2\pi)^{3}}{\rm Li}_{3}(e^{-2\pi(T-U)})+O\left(e^{-2\pi S}\right)\bigg{)}\,,\end{split}$ (3.17)
where the coefficients $c_{1}(n)$ are defined via
$\begin{split}\frac{E_{4}E_{6}}{\eta^{24}}(q)=\sum_{n=-1}^{\infty}c_{1}(n)=\frac{1}{q}-240-141444q+\ldots\,.\end{split}$ (3.18)
The cubic part is
$\begin{split}\mathcal{F}_{0}^{\rm het,cubic}=STU+\frac{1}{3}U^{3}+T^{2}U\,.\end{split}$ (3.19)
Note that only the first term is at tree-level in $g_{s}^{\rm het}$, whereas the two $S$ independent ones and all remaining $S$ independent ones from (3.17) make the full one-loop correction. The remaining unknown terms come from heterotic $NS5$-brane instantons. One can check that the type IIA and heterotic cubic terms directly match reflecting the duality $\mathcal{F}_{0}^{\rm IIA}=\mathcal{F}_{0}^{\rm het}$.
Moreover, exploiting this duality the Gopakumar-Vafa invariants in the class
$\begin{split}C=i(kT+lU)=(k+l)t_{1}+kt_{3}\end{split}$ (3.20)
can be directly read-off from (3.17)
$\begin{split}\alpha_{0}^{(k+l,0,k)}=-2c_{1}(kl)\qquad{\rm for}\quad k,l\geq 0\,\ {\rm and}\ (k,l)=(1,-1)\,.\end{split}$ (3.21)
Note that $l$ can also take negative values without violating the Kähler cone condition. Hence, we are in a situation where we happen to know the subset of GV invariants for $D2$-branes not wrapping the large base explicitly in terms of a modular form. Note that this implies that these GV invariants grow like $c_{1}(n)\sim\exp(4\pi\sqrt{n})$ for large $n$, which is correlated with the emergent string limit that we are taking.
3.2.1 Yukawa couplings from Schwinger integrals
All these preliminary considerations were in some weakly coupled limit. Now we analyse how the computation could be carried out in the co-scaled infinite distance limit $\tau_{2}\to\lambda\tau_{2}$, $g^{\rm IIA}_{s}\to\sqrt{\lambda}g^{\rm IIA}_{s}$, where the 4D Planck scale remains constant. In this limit, the light towers of states with a mass scale below the species scale $\Lambda_{\rm sp}\sim M_{\rm pl}/\sqrt{\lambda}$ are strings arsing from $NS5$-branes wrapping the $K3$-fiber and $D2$-$D0$ bound states wrapping 2-cycles of the $K3$. The contribution of these $D2$-$D0$ bound states to ${\cal F}_{0}$ is given by the Schwinger integral (2.2).
To be concrete, let us discuss the zero point Yukawa-coupling $Y^{(0)}_{UUU}$. To evaluate it, one first needs to determine all GV invariants $\alpha_{0}^{(n_{1},0,n_{3})}$ for 1/2 BPS $D2$-branes wrapping the 2-cycles in the $K3$-fiber. In general this is a horrendous problem that only a “CY-demon” can solve by pure counting. However, as just explained, we are in the fortunate situation that we know these GV invariants already from computations in the weakly coupled regime. Hence, the “CY-demon” will eventually find the same numbers. Therefore, for the expression for the Yukawa coupling (2.4) the demon will write
$\begin{split}Y_{UUU}=Y_{UUU}^{(0)}+2\cdot 240\sum_{l=1}^{\infty}l^{3}\,{q_{2}^{l}\over(1-q_{2}^{l})}-\sum_{k,l=1}^{\infty}2c_{1}(kl)\,l^{3}{q_{1}^{k}q_{2}^{l}\over(1-q_{1}^{k}q_{2}^{l})}-2\frac{q_{1}}{q_{1}-q_{2}}\,,\end{split}$ (3.22)
where $q_{1}=\exp(-2\pi T)$, $q_{2}=\exp(-2\pi U)$ and we have used $c_{1}(-1)=1$ and $c_{1}(0)=-240$. One realizes that the second term on the right hand side can be expressed via the Eisenstein series
$\begin{split}E_{4}(q)=1+240\sum_{n=1}^{\infty}\frac{n^{3}q^{n}}{1-q^{n}}\,\end{split}$ (3.23)
as $2(E_{4}(q_{2})-1)$. According to (2.4), the diverging zero point Yukawa coupling is given by
$\begin{split}Y^{(0)}_{UUU}&=240\sum_{l=1}^{\infty}l^{3}-\sum_{k,l=1}^{\infty}c_{1}(kl)\,l^{3}-c_{1}(-1)\cdot(-1)^{3}\\[2.84544pt] &=-2\cdot 240\sum_{l,d=1}^{\infty}l^{3}+2\sum_{k,l,d=1}^{\infty}c_{1}(kl)\,l^{3}+2\sum_{d=1}^{\infty}c_{1}(-1)\cdot(-1)^{3}\,,\end{split}$ (3.24)
where, as for the example in section 2.2, we introduced the sum $\sum_{d=1}^{\infty}1=\zeta(0)=-1/2$. Following the same example, we can now regularize this expression by introducing dummy variables $p_{1},p_{2}<1$ with $p_{1}<p_{2}$ and eventually take the limit $p_{1},p_{2}\to 1$. Thus, we get
$\begin{split}Y^{(0)}_{UUU}\Big{|}_{\rm reg}:=&\lim_{p_{1},p_{2}\to 1}\left[-2\cdot 240\sum_{l,d=1}^{\infty}l^{3}\,p_{2}^{ld}+2\sum_{k,l,d=1}^{\infty}c_{1}(kl)\,l^{3}\,p_{1}^{kd}\,p_{2}^{ld}-2\sum_{d=1}^{\infty}p_{1}^{d}\;p_{2}^{-d}\,\right]\\[2.84544pt] =&\lim_{p_{1},p_{2}\to 1}\left[2-2E_{4}(p_{2})+2\sum_{k,l=1}^{\infty}c_{1}(kl)l^{3}{p_{1}^{k}\,p_{2}^{l}\over(1-p_{1}^{k}\,p_{2}^{l})}+2{p_{1}\over(p_{1}-p_{2})}\right]\,.\end{split}$ (3.25)
Note that the geometric series on the far right-hand side will only be convergent if $p_{1}<p_{2}$. Clearly, after these steps the regularized expression has the same form as the non zero-point Yukawa coupling in (3.22) but it is important to note that we are not taking the $p_{1},p_{2}\to 0$ limit, where this expression nicely converges but the opposite $p_{1},p_{2}\to 1$ limit. In order to perform this limit and to isolate the singularity, we employ a useful mathematical relation, the so-called Harvey-Moore identity [42]
$\begin{split}-E_{4}(p_{2})+\sum_{k,l,d=1}^{\infty}c_{1}(kl)\,l^{3}\,p_{1}^{kd}p_{2}^{ld}+\frac{p_{1}}{p_{1}-p_{2}}=\frac{E_{4}(p_{1})E_{6}(p_{1})}{\eta^{24}(p_{1})}\frac{E_{4}(p_{2})}{j(p_{2})-j(p_{1})}\,,\end{split}$ (3.26)
where the right hand side is a modular form of bi-degree $(-2,4)$.
As for the pedagogical toy example in section 2.2, we now write $p_{1}=\exp(-2\pi/\Lambda_{1})$, $p_{2}=\exp(-2\pi/\Lambda_{2})$ with $\Lambda_{1}<\Lambda_{2}$ and apply a modular $S$-transformation to the right-hand side of (3.26). In this way, we get
$\begin{split}Y_{UUU}^{(0)}&=\lim_{\Lambda_{1},\Lambda_{2}\rightarrow\infty}\left(2-2{\Lambda_{2}^{4}\over\Lambda_{1}^{2}}\frac{E_{4}(i\Lambda_{1})E_{6}(i\Lambda_{1})}{\eta^{24}(i\Lambda_{1})}\frac{E_{4}(i\Lambda_{2})}{j(i\Lambda_{2})-j(i\Lambda_{1})}-{\rm Div}\right)\\ &=\lim_{\Lambda_{1},\Lambda_{2}\rightarrow\infty}\left(2-2{\Lambda_{2}^{4}\over\Lambda_{1}^{2}}\Big{(}1+O(e^{-2\pi(\Lambda_{2}-\Lambda_{1})})\Big{)}-{\rm Div}\right)\,.\end{split}$ (3.27)
Here we realize the compelling pattern that in the limit, the second term in (3.27) contains only a diverging term and exponentially suppressed terms. Hence, after minimally subtracting the divergent ${\Lambda_{2}^{4}/\Lambda_{1}^{2}}$ term and taking the limit we arrive at the final result
$\begin{split}Y_{UUU}^{(0)}=2\,,\end{split}$ (3.28)
which is indeed the cubic term in the one-loop heterotic Yukawa coupling (3.19). Recall that this corresponds to a tree-level TIN in the type IIA dual model.
Thus, what we have effectively done here is the following: since we do not have a “CY demon”, we used the knowledge of the GV invariants derived in the weakly coupled region from the dual heterotic string (or as we will see later from type IIA mirror symmetry) to define the regularized value of the zero point Yukawa coupling, arising from the Schwinger integral over $D2$-$D0$ bound states (in the strongly coupled region), as
$\begin{split}Y_{UUU}^{(0)}\vcentcolon=-\lim_{U,T\rightarrow 0}\left[\big{(}i\,\partial_{U}^{3}\mathcal{F}_{0}^{\rm het}-\kappa_{UUU}\big{)}-{\rm Div}\right]\,,\end{split}$ (3.29)
where Div denotes all terms that diverge in the limit. We note that the cubic term $\kappa_{UUU}$ can be obtained from the opposite limit
$\begin{split}\kappa_{UUU}=\lim_{U,T\rightarrow\infty}i\,\partial_{U}^{3}\mathcal{F}_{0}^{\rm het}\,.\end{split}$ (3.30)
The definition (3.29) means that one gets precisely $Y_{UUU}^{(0)}=\kappa_{UUU}$, if the $U,T\rightarrow 0$ limit of $\partial_{U}^{3}\mathcal{F}_{0}^{\rm het}$ itself only gives divergent and vanishing terms, i.e. no further constant terms. Thus, it is satisfactory that the proposed regularization involves the TIN $\kappa$ in an explicit way allowing us to formulate the previous simple criterion.
One could now use this definiton to evaluate the remaining three zero point Yukawa couplings $Y_{UUT}^{(0)}$, $Y_{UTT}^{(0)}$ and $Y_{TTT}^{(0)}$. Instead of doing this in detail, we just mention that $Y_{TTT}^{(0)}=0$ can be obtained from an analogous computation in the chamber ${\rm Re}(U)>{\rm Re}(T)$ and an analytic continuation of the ${\rm Li}_{3}(e^{-2\pi(T-U)})$-term in (3.17). The couplings $Y_{UUT}^{(0)}$ and $Y_{UTT}^{(0)}$ will require a different technique, as the Harvey-Moore formula cannot be applied directly.
To summarize, for an emergent string limit we have successfully invoked a regularization of the zero point Yukawa couplings so that the Schwinger integration over the light towers of $D2$-$D0$ bound states at strong coupling reproduced a subset of the classical (tree-level) TINs. Even though we are not yet taking the isotropic M-theory limit, this computation serves as proof of principle of how this regularization can be defined. Of course we are not yet getting full emergence as the term $\kappa_{STU}$ is at (heterotic $NS5$-brane) string tree-level and not captured by the one-loop Schwinger integral. The next step is to move forward to the isotropic M-theory limit, where we expect the full weakly coupled type IIA tree-level prepotential to be contained in the Schwinger one-loop integral at strong coupling.
4 Isotropic M-theory limit of CYs
Now we want to generalize these computations to the ultimate case, namely to isotropic M-theory limits where the full homology lattice contributes and where the GV invariants generically scale exponentially as $\exp(\gamma\beta)$.
Let us note that the example from the previous section shares one essential feature with the resolved conifold, namely that the 2-cycles wrapped by the light $D2$-branes are allowed to shrink to zero size (in string units). As is obvious from our construction, this feature was crucial in defining the regularization of (2.5) via the limit $t\to 0$ and employing the modular properties of the Yukawa couplings to isolate the divergence and subtract it. However, as already observed in the seminal work [28] and briefly reviewed in appendix A, on the Quintic threefold, the exponential degeneracy of the GV invariants is correlated with the appearance of a divergence of the prepotential not at zero but at a finite value $t=t_{c}$ set by the location of the conifold singularity. Note that this is not the minimal value of ${\rm Im}(t)$ which is at the Landau-Ginzburg point. Indeed, the Yukawa-couplings are regular at that point so that it does not seem reasonable to regularize our infinite sums by taking the limit to this point.
As shown in appendix B, the conifold point is nothing else than the quantum corrected singular point in the middle of the classical phase diagram of a GLSM. Thus, in a certain sense we are indeed taking a limit towards zero size, namely towards vanishing FI parameter $\xi=0$. Since due to quantum effects this is not really possible, the most natural thing to do is to take the limit towards its quantum corrected value.
4.1 Emergence of TINs on CYs with $h_{11}=1$
For concreteness, let us now consider CYs with one Kähler modulus and therefore no fibration structure. In this case, there is no obvious toroidal structure present and one does not expect the Yukawa coupling to contain any modular forms. As mentioned, the Yukawa couplings in the large radius chart still have a divergence of the form
$\begin{split}Y_{ttt}\sim{1\over(t-t_{c})\log^{2}(t-t_{c})}+\ldots\,,\end{split}$ (4.1)
which is not at zero but at the location of the conifold point $t=t_{c}$. The same holds for the total list of 14 CYs with $h_{11}=1$ that we provide here for convenience (see e.g. [26])
$\begin{split}&\mathbb{P}_{1^{4},4,6}[12,2]_{2.302}\,,\ \mathbb{P}_{1^{3},2,5}[10]_{2.099}\,,\ \mathbb{P}_{1^{2},2^{2},3^{2}}[6,6]_{1.874}\,,\ \mathbb{P}_{1^{4},4}[8]_{1.695}\,,\ \mathbb{P}_{1^{3},2^{2},3}[6,4]_{1.563}\,,\\[2.84544pt] &\mathbb{P}_{1^{4},2}[6]_{1.421}\,,\mathbb{P}_{1^{5},3}[6,2]_{1.334}\,,\ \mathbb{P}_{1^{4},2^{2}}[4,4]_{1.255}\,,\ \mathbb{P}_{1^{5}}[5]_{1.208}\,,\ \mathbb{P}_{1^{5},2}[4,3]_{1.115}\,,\\[2.84544pt] &\mathbb{P}_{1^{5}}[4,2]_{1.029}\,,\ \mathbb{P}_{1^{6}}[3,3]_{0.975}\,,\ \mathbb{P}_{1^{7}}[3,2,2]_{0.891}\,,\ \mathbb{P}_{1^{8}}[2,2,2,2]_{0.807}\,,\end{split}$ (4.2)
where the lower index is the value of ${\rm Im}(t_{c})$ at the conifold singularity. In fact, in [28] this divergence was utilized to derive the asymptotic behavior of the GV invariants
$\begin{split}\alpha_{0}^{\beta}\sim{e^{2\pi\beta\,{\rm Im}(t_{c})}\over\beta^{3}\log(\beta)^{2}}\,\,,\end{split}$ (4.3)
which is briefly reviewed in appendix A. Crucially, for defining a regularization of the zero point Yukawa couplings, the divergence at the conifold singularity encodes information about the exponential degeneracy of the GV invariants. Therefore, it does not make sense to define the regularization of the zero point Yukawa coupling via a $t\to 0$ limit, but rather as
$\begin{split}Y_{ttt}^{(0)}&:=-\lim_{t\to t_{c}}\left[\sum_{n=1}^{\infty}\alpha_{0}^{n}\,n^{3}{q^{n}\over 1-q^{n}}-{\rm Div}\right]=-\lim_{t\to t_{c}}\left[\Big{(}\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}-\kappa_{ttt}\Big{)}-{\rm Div}\right]\end{split}$ (4.4)
with $q=\exp(2\pi it)$ and having set $(2\pi i)^{3}/g_{s}^{2}=1$. As usual, before taking the limit to the conifold point, we minimally subtract the divergence for $(t-t_{c})\to 0$. To explain the logic, recall that the computation is now done in the strong coupling M-theory limit, where we integrate out the full tower of 1/2-BPS $D2$-$D0$ bound states. Of course, only a “CY demon” can do this in practise, but we are lucky in that we happen to know the exact result from the weakly coupled regime via non-renormalization theorems and mirror symmetry. Thus, the result that the “CY-demon” obtains after carrying out this infinite sum should agree with the weakly coupled result in the large complex structure regime $\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}$ minus the triple intersection number $\kappa_{ttt}$, that is not contained in the sum $\sum_{n=1}^{\infty}\alpha_{0}^{n}\,n^{3}{q^{n}/(1-q^{n})}$.
Now, we have arrived at an expression that is very similar to the other regularized expressions (2.13) and (3.29) encountered so far, where we were dealing with modular forms. Instead of applying a modular S-transformation, next we have to evaluate $\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}$ around the conifold point and disentangle the divergent, the constant and the suppressed contributions in the $(t-t_{c})\to 0$ limit. We have succeeded in defining a working regularization, if the following regularization condition holds:
The $t\rightarrow t_{c}$ limit of $\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}$ only gives divergent and vanishing terms, i.e. no further constant terms. Then (4.4) yields precisely $Y_{ttt}^{(0)}=\kappa_{ttt}$.
4.1.1 Example: Quintic
Let us consider the Quintic, i.e. $\mathbb{P}_{4}[5]$, as a concrete example, where we can rely on well known results. The weakly coupled Yukawa coupling $\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}$ was already computed in the seminal paper [28] using mirror symmetry. Here we are just collecting a couple of useful relations obtained in [28] and also in [43, 44] and refer the reader to the original literature. The mirror of the Quintic has one complex structure modulus, which is given by the deformation of the degree five hypersurface constraint
$\begin{split}\sum_{i=1}^{5}z_{i}^{5}-(5\psi)\,z_{1}z_{2}z_{3}z_{5}z_{5}=0\end{split}$ (4.5)
in $\mathbb{P}^{4}$. The conifold singularity is located at $\psi=1$. A basis of local periods can be determined as solutions to the Picard-Fuchs equation and in the large complex structure regime can be expressed as
$\begin{split}\pi_{{\rm LCS},1}&=\omega_{0}(z)=\sum_{n=0}^{\infty}{(5n)!\,z^{n}\over(n!)^{5}}\,\\ \pi_{{\rm LCS},2}&=\omega_{0}\log z+z\,\sigma_{1}(z)\\ \pi_{{\rm LCS},3}&=\omega_{0}\log^{2}z+2z\,\sigma_{1}(z)\log z+z\,\sigma_{2}(z)\\ \pi_{{\rm LCS},4}&=\omega_{0}\log^{3}z+3z\,\sigma_{1}(z)\log^{2}z+3z\,\sigma_{2}(z)\log z+z\,\sigma_{3}(z)\end{split}$ (4.6)
with $z=1/(5\psi)^{5}$ and $|z|<5^{-5}$. The $\sigma_{i}(z)=\sum_{n=0}^{\infty}a_{i,n}z^{n}$ are infinite series in $z$, starting with a constant. Their precise form will not be important for our purposes. The next step is to map these periods to to an integral symplectic basis, which allows to read off the mirror map and the prepotential via
$\begin{split}\Pi_{\rm LCS}=\left(\begin{matrix}X_{0}\\ X_{1}\\ F_{1}\\ F_{0}\end{matrix}\right)=\omega_{0}\left(\begin{matrix}1\\ t\\ \partial_{t}{\cal F}_{0}\\ 2{\cal F}_{0}-t\partial_{t}{\cal F}_{0}\end{matrix}\right)\,.\end{split}$ (4.7)
Here $t=iT$ denotes the complexified Kähler modulus of the Quintic that enjoys an expansion
$\begin{split}t={X_{1}\over X_{0}}={1\over 2\pi i}\log z+O(z)\,.\end{split}$ (4.8)
Inverting this mirror map one can express the prepotential around the large complex structure point as
$\begin{split}{\cal F}_{0}={5\over 6}t^{3}+{25i\zeta(3)\over 2\pi^{3}}+{\rm instantons}\,.\end{split}$ (4.9)
The radius of convergence of the large complex structure patch is precisely $|\psi|=1$ so that the conifold singularity lies on the boundary. Therefore, for determining the limit of the Yukawa coupling (4.4) this patch is not convenient and one needs to know the periods in the conifold patch and how the two patches are correctly glued together. For the quintic this problem has been solved e.g. in [44] and here we just recall their results up to the level of detail needed for our purposes.
In the conifold patch, it is convenient to introduce the coordinate $u=1-\psi^{-5}$ which vanishes at the conifold locus. Then one can solve the Picard-Fuchs equation in this coordinate, obtain a basic of periods and finally determine a transition matrix to an integral symplectic basis $\Pi_{\rm C}$, which on the overlap with the LCS patch matches the former basis $\Pi_{\rm LCS}$. This integral symplectic basis takes the form
$\begin{split}X_{0}&=-\rho_{1}(u){u\over 2\pi i}\log u+\rho_{2}(u)\,\\ X_{1}&=\rho_{3}(u)\,\\ F_{1}&=\rho_{4}(u)\,\\ F_{0}&=u\,\rho_{1}(u)\,,\end{split}$ (4.10)
where the $\rho_{i}(u)=\sum_{n=0}^{\infty}b_{i,n}u^{n}$ are infinite series in $u$. Note that the logarithmic term in $X_{0}$ leads to the monodromy $X_{0}\to X_{0}-F_{0}$ around the conifold.
Now we can read off the Kähler $t$ in terms of the conifold coordinate $u$, via
$\begin{split}t={X_{1}\over X_{0}}={b_{4,0}\over b_{3,0}}\sum_{m=0}^{\infty}(u\log u)^{m}\,c_{m}(u)\,,\end{split}$ (4.11)
where $c_{m}(u)$ denote infinite series in $u$ with $c_{0,0}=1$. Using the concrete values $b_{4,0}=i1.29357...$, $b_{3,0}=1.07073...$ we obtain in the $u\to 0$ limit $t_{c}=iT_{c}=i1.20812...$. Note that at leading order we get
$\begin{split}t-t_{c}=t_{c}\,c_{1,0}\,u\log u+\ldots\,.\end{split}$ (4.12)
Next we recall that the Yukawa coupling can be expressed as [28]
$\begin{split}\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}={1\over\omega_{0}^{2}}\kappa_{\psi\psi\psi}{1\over(dt/d\psi)^{3}}\end{split}$ (4.13)
with
$\begin{split}\kappa_{\psi\psi\psi}=\left(\frac{2\pi i}{5}\right)^{3}{5\psi^{2}\over(1-\psi^{5})}\,.\end{split}$ (4.14)
Using $u=1-\psi^{-5}$ and the general structure of the periods at the conifold, one realizes that the building blocks in (4.13) can be expressed as
$\begin{split}{dt\over du}&=\log u\left({d_{-1}(u)\over\log u}+\sum_{n=0}^{\infty}(u\log u)^{n}\,d_{n}(u)\right)\\ {1\over\omega_{0}^{2}}&=\sum_{n=0}^{\infty}(u\log u)^{n}\,e_{n}(u)\\ {\kappa_{\psi\psi\psi}\over(\partial_{\psi}u)^{3}}&=-{1\over 25}{1\over u(1-u)^{3}}\,,\end{split}$ (4.15)
where the $d_{n}(u)$ and $e_{n}(u)$ are again infinite series in $u$. Putting everything together and expanding for small $u$, we find the general expansion
$\begin{split}\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}={1\over u\log^{3}u}\left(\sum_{n,k=0}^{\infty}{(u\log u)^{n}\over\log^{k}u}a_{n,k}(u)\right)\,.\end{split}$ (4.16)
Consistent with [28], one sees that the leading divergence is
$\begin{split}\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}\sim{1\over u\log^{3}u}+\ldots\sim{1\over(t-t_{c})\log^{2}(t-t_{c})}+\ldots\,,\end{split}$ (4.17)
where we used (4.12). However, for $n=0$ all the constant terms in $a_{0,k}(u)=a_{0,k,0}+O(u)$ lead to singular terms
$\begin{split}\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}=\sum_{k=0}^{\infty}{a_{0,k,0}\over u\log^{k+3}u}+{\rm Reg}\,,\end{split}$ (4.18)
which according to our philosophy are minimally subtracted. With all the remaining terms being regular, the final and essential question is whether they all go to zero or whether there remains a constant contribution. That this is not the case can be straightforwardly seen by noting that the cancellation of the $\log u$ term requires $n=k+3$, for which still a term $u^{k+2}a_{k+3,k}(u)$ remains that goes to zero for $u\to 0$. Hence there is no constant term and our proposed regularization (4.4) indeed yields
$\begin{split}Y_{ttt}^{(0)}&=-\lim_{t\to t_{c}}\left[\Big{(}\partial^{3}_{t}{\cal F}_{0}\big{|}_{\rm weak}-\kappa_{ttt}\Big{)}-{\rm Div}\right]=\kappa_{ttt}=5\,.\end{split}$ (4.19)
We expect the computation to be completely analogous for the list of 14 CYs with $h_{11}=1$ presented in (LABEL:listCYh11).
Having a closer look at the computation one realizes that the avoidance of a constant contribution in the limit $t\to t_{c}$ can be traced back to the appearance of the $u\log u$ term in the period $X_{0}$, i.e. that $\log u$ is always accompanied by a linear factor $u$. This is of course a consequence of the monodromy around the conifold point.
It could be that our findings are just an artifact of the too simple choice of CY manifolds with one Kähler modulus. Therefore, next we generalize the computation to CYs with two Kähler moduli. Moreover, we also analyze whether the regularization condition is only true for approaching conifold singularities or whether it also holds for more general degeneration limits. For $h_{11}=2$ it also happens that the codimension one degeneration loci intersect in more than one point so that the question arises which point to choose and also along which family of path it should be approached in taking the limit. To approach these questions, in the following sections we consider the two CYs $\mathbb{P}_{1,1,1,6,9}[18]$ and $\mathbb{P}_{1,1,2,2,6}[12]$.
4.2 Emergence of TINs on $\mathbb{P}_{1,1,1,6,9}[18]$
We start with the elliptically fibered CY manifold from section 3.1, namely $\mathbb{P}_{1,1,1,6,9}[18]$ with Hodge numbers $(h_{11},h_{21})=(2,272)$. The two complex structure moduli of the mirror dual CY are parametrized by the deformations of the hypersurface constraint
$\begin{split}p=z_{1}^{18}+z_{2}^{18}+z_{3}^{18}+z_{4}^{3}+z_{5}^{2}-18\psi\,z_{1}z_{2}z_{3}z_{4}z_{5}-3\phi\,z_{1}^{6}z_{2}^{6}z_{3}^{6}\,.\end{split}$ (4.20)
Introducing the combinations
$\begin{split}\overline{x}=-\frac{1}{2^{2}3^{8}}{\phi\over\psi^{6}}\,,\qquad\overline{y}=-{1\over\phi^{3}}\,,\end{split}$ (4.21)
the two conifold degeneration loci of the manifold are
$\begin{split}\Delta_{1}&=(1-\overline{x})^{3}-\overline{x}^{3}\overline{y}=0\\ \Delta_{2}&=1+\overline{y}=0\,.\end{split}$ (4.22)
The LCS point is at $\overline{x}=\overline{y}=0$ and one can solve the corresponding Picard-Fuchs equations in the vicinity of this point, determine an integral symplectic bases of periods $\Pi^{\rm LCS}_{I}$, $I=1,\ldots,6$, such that the mirror map is
$\begin{split}t_{1}={\Pi^{\rm LCS}_{2}\over\Pi^{\rm LCS}_{1}}={1\over 2\pi i}\log(\overline{x})+\ldots\,,\qquad t_{2}={\Pi^{\rm LCS}_{3}\over\Pi^{\rm LCS}_{1}}={1\over 2\pi i}\log(\overline{y})+\ldots\,.\end{split}$ (4.23)
Here $t_{1}$ measures the size of the toroidal fiber and $t_{2}$ the size of the curve $\mathbb{P}^{1}\subset\mathbb{P}^{2}$. Inverting the mirror map, the prepotential comes out as
$\begin{split}{\cal F}_{0}|_{\rm weak}={3\over 2}t_{1}^{3}+{3\over 2}t_{1}^{2}t_{2}+{1\over 2}t_{1}t^{2}_{2}+{135i\zeta(3)\over 4\pi^{3}}+{\rm instantons}\,.\end{split}$ (4.24)
We are particularly interested in the Yukawa couplings, for which the exact B-side Yukawas were provided in [45, 33], which in terms of $(\bar{x},\bar{y})$ read [30]
$\begin{split}\kappa_{\bar{x}\bar{x}\bar{x}}&=\frac{9i}{8\pi^{3}}\frac{1}{\bar{x}^{3}((1-\bar{x})^{3}-\bar{x}^{3}\bar{y})}\,,\quad\ \kappa_{\bar{x}\bar{x}\bar{y}}=\frac{i}{8\pi^{3}}\frac{3(1-\bar{x})}{\bar{x}^{2}\bar{y}((1-\bar{x})^{3}-\bar{x}^{3}\bar{y})}\,,\\ \kappa_{\bar{x}\bar{y}\bar{y}}&=\frac{i}{8\pi^{3}}\frac{(1-\bar{x})^{2}}{\bar{x}\bar{y}^{2}((1-\bar{x})^{3}-\bar{x}^{3}\bar{y})}\,,\quad\kappa_{\bar{y}\bar{y}\bar{y}}=\frac{i}{24\pi^{3}}\frac{1-3\bar{x}+3\bar{x}^{2}}{\bar{y}^{2}(1+\bar{y})((1-\bar{x})^{3}-\bar{x}^{3}\bar{y})}\,.\end{split}$ (4.25)
Note that these expressions are valid throughout the entire complex structure moduli space.
For regularizing the zero point Yukawa couplings (1.1) we now want to take the limit to a point in the complex structure moduli space, where the CY develops a singularity, which intuitively encodes the asymptotic growth of the GV invariants in all directions of the homology lattice $H_{2}(X,\mathbb{Z})$. The natural choice are the intersection points of the two conifold loci (4.22). According to [33], there exist three such points, where the two points $\overline{x}_{0}=(3\pm\sqrt{3}i)/6$ and $\overline{y}_{0}=-1$ are obvious and the third one at $\overline{w}_{0}=1/\overline{x}_{0}=0$, $\overline{y}_{0}=-1$ is actually a triple intersection with another orbifold singularity $D_{0}$ at $\psi=0$. This is shown on the left in figure 1, which is essentially taken from [33].
[…]
Figure 1: Schematic view of the intersections of degeneration loci $D_{1}=\{\Delta_{1}=0\}$, $D_{2}=\{\Delta_{2}=0\}$, $D_{0}$ and their resolution.
To get a smooth moduli space with normal crossings of the degeneration loci, one needs to perform a resolution of the triple intersection. This is done via a blow-up, which introduces a divisor $E_{3}$ which has normal crossings with the three degeneration loci as shown on the right in figure 1.
4.2.1 Emergence at the point $P_{+}$
First, let us consider the intersection point $P_{+}$ with $\overline{x}_{0}=(3+\sqrt{3}i)/6$ and $\overline{y}_{0}=-1$, where as shown in [46] it is convenient to introduce the local coordinates
$\begin{split}x_{1}=1-{\overline{x}\over\overline{x}_{0}}\,,\qquad\quad x_{2}=\left(1-{\overline{x}\over\overline{x}_{0}}\right)^{-1}\left(1-{\overline{y}\over\overline{y}_{0}}\right)\,\end{split}$ (4.26)
for which the two conifold degeneration loci become
$\begin{split}\Delta_{1}={\textstyle{(1-i\sqrt{3})\over 2}}x_{1}\big{(}1+O(x_{i})\big{)}\,,\qquad\quad\Delta_{2}=x_{1}x_{2}\,.\end{split}$ (4.27)
In appendix B, we show that the intersection of the two conifold loci (4.22) corresponds to the quantum corrected singular point in the middle of the classical phase diagram (for the Fayet-Iliopoulos terms) of the GLSM. Note that $(x_{1},x_{2})$ should be regarded rather as (complex) polar coordinates around $P_{+}$, where $x_{1}$ is the radial direction and $x_{2}$ related to the angle, which is evident from
$\begin{split}{\overline{y}-\overline{y}_{0}\over\overline{x}-\overline{x}_{0}}={\overline{y}_{0}\over\overline{x}_{0}}x_{2}=\tan\phi\end{split}$ (4.28)
and that $P_{+}$ is reached for $x_{1}=0$ independent of $x_{2}$. Like for radial coordinates this means that the map is not invertible at $P_{+}$.
Then, our proposal is that the TINs $\kappa_{t_{i}t_{j}t_{k}}$ are given by the regularization
$\begin{split}Y_{t_{i}t_{j}t_{k}}^{(0)}=-\lim\limits_{\begin{subarray}{c}t_{1}\to t_{1,c}\\ t_{2}\to t_{2,c}\end{subarray}}\left[\Big{(}\partial_{t_{i}}\partial_{t_{j}}\partial_{t_{k}}{\cal F}_{0}\big{|}_{\rm weak}-\kappa_{t_{i}t_{j}t_{k}}\Big{)}-{\rm Div}\right]=\kappa_{t_{i}t_{j}t_{k}}\,.\end{split}$ (4.29)
Therefore, to actually evaluate the expression (4.29) we need to know the third derivatives of the prepotential in a vicinity of the intersection point $P_{+}$ of the two conifold loci. For that purpose, we first solve the Picard-Fuchs equations in local coordinates around the point $(x_{1},x_{2})=(0,0)$ and then determine an integral symplectic basis which is required to be the continuation of the integral symplectic basis around the LCS point $(\overline{x},\overline{y})=(0,0)$. That means that on the overlap of two local charts around these points, they have to coincide. To determine the numerical transition matrices with high enough precision, we have computed all period sets up to total order $40$ in the respective local variables $(x_{1},x_{2})$.
Similar computations have been performed in [15, 47] and for our purposes we have been following this general recipe to arrive at the following results [^5] We are grateful to Rafael Álvarez-García who helped us learning how to perform these computations., whose detailed derivation will be reported in [27]. Eventually, the periods around the point $(x_{1},x_{2})=(0,0)$ take the schematic form
$\begin{split}X_{\rm 0}&=x_{1}P_{0}\log(x_{1})+Q_{0}\,\\[2.84544pt] X_{\rm 1}&=x_{1}x_{2}P_{1}\log(x_{1}x_{2})+Q_{1}\,\\[2.84544pt] X_{\rm 2}&=x_{1}x_{2}P_{2}\log(x_{1}x_{2})+x_{1}\tilde{P}_{2}\log(x_{1})+Q_{2}\,\\[2.84544pt] F_{\rm 2}&=x_{1}x_{2}\hat{P}_{2}\log(x_{1}x_{2})+\hat{Q}_{2}\,\\[2.84544pt] F_{\rm 1}&=x_{1}\hat{P}_{1}\log(x_{1})+\hat{Q}_{1}\,\\[2.84544pt] F_{\rm 0}&=x_{1}\hat{P}_{0}\log(x_{1})+\hat{Q}_{0}\,,\\ \end{split}$ (4.30)
where the $P_{i},Q_{i},\hat{P}_{i},\hat{Q}_{i}$ are polynomials in the coordinates $x_{1},x_{2}$. The periods clearly feature the expected logarithmic terms at the locations of the two conifold singularities (4.27).
Then, the two Kähler moduli enjoy the expansions
$\begin{split}t_{i}=\sum_{n=0}^{\infty}\big{(}x_{1}\log x_{1}\big{)}^{n}a_{n}^{(i)}+(x_{1}x_{2})\log(x_{1}x_{2})\sum_{n=0}^{\infty}\big{(}x_{1}\log x_{1}\big{)}^{n}\,b_{n}^{(i)}\qquad i=1,2\,,\end{split}$ (4.31)
where $a_{n}^{(i)},b_{n}^{(i)}$ are polynomials in $x_{1},x_{2}$, where however an $x_{2}^{n}$ term is always accompanied by at least one factor of $x_{1}$. In the limit $x_{1}=0$, the values of $t_{i}$ are determined by the constant terms in $a_{0}^{(i)}$ which give
$\begin{split}t^{(0)}_{1}\approx 0.137+0.991i\,,\qquad\qquad t^{(0)}_{2}\approx 0.545+0.427i.\end{split}$ (4.32)
In figure 2 we show a parametric plot of $({\rm Im}(t_{1}),{\rm Im}(t_{2}))$ around the singularity at $(t^{(0)}_{1},t^{(0)}_{2})$. This shows that also in the Kähler moduli space the singularity can be approached from various directions.
Figure 2: Parametric plot of $({\rm Im}(t_{1}-t_{1}^{(0)}),{\rm Im}(t_{2}-t_{2}^{(0)}))$ for varying $(x_{1},x_{2})$ with ${\rm Im}(x_{i})=0$ and $0\leq{\rm Re}(x_{1})\leq 0.1$, $0.01\leq{\rm Re}(x_{2})\leq 2.0$.
Now we are ready to compute the Yukawa couplings (4.25). For that purpose we first transform them to the new coordinates $(x_{1},x_{2})$ and then to the Kähler moduli $(t_{1},t_{2})$ via
$\begin{split}\partial_{t_{i}}\partial_{t_{j}}\partial_{t_{k}}\mathcal{F}_{0}|_{\rm weak}=\frac{1}{\omega_{0}^{2}}\sum_{\alpha,\beta,\gamma}\frac{\partial x_{\alpha}(t)}{\partial t_{i}}\frac{\partial x_{\beta}(t)}{\partial t_{j}}\frac{\partial x_{\gamma}(t)}{\partial t_{k}}\kappa_{x_{\alpha}x_{\beta}x_{\gamma}}\,.\end{split}$ (4.33)
Using (4.25), for the $\kappa_{x_{\alpha}x_{\beta}x_{\gamma}}$ we find the behavior
$\begin{split}\kappa_{x_{1}x_{1}x_{1}}&={1\over x_{1}}P_{1}(x_{1},x_{2})\,,\qquad\kappa_{x_{1}x_{1}x_{2}}=P_{2}(x_{1},x_{2})\,,\\ \kappa_{x_{1}x_{2}x_{2}}&=x_{1}^{2}P_{3}(x_{1},x_{2})\,,\qquad\;\kappa_{x_{2}x_{2}x_{2}}={x_{1}^{2}\over x_{2}}P_{4}(x_{1},x_{2})\,,\end{split}$ (4.34)
where the $P_{i}(x_{1},x_{2})$ are polynomials in $(x_{1},x_{2})$ having a non-vanishing constant term.
For computing the partial derivatives $\partial x_{\alpha}(t)/\partial t_{j}$ one needs to the invert the mirror map, which for the periods (4.30) does not allow for an iterated procedure. We circumvent this problem, by assuming that the $x_{i}$ are non-vanishing and applying the inverse function theorem for holomorphic functions to determine the derivatives $\partial x_{\alpha}/\partial t_{j}$ implicitly. To do so, we need to invert the Jacobian of the mirror map
$\begin{split}J_{\rm mir}(x_{1},x_{2})=\begin{pmatrix}\frac{\partial t_{1}}{\partial x_{1}}&\frac{\partial t_{1}}{\partial x_{2}}\\ \frac{\partial t_{2}}{\partial x_{1}}&\frac{\partial t_{2}}{\partial x_{2}}\\ \end{pmatrix}\,,\end{split}$ (4.35)
allowing us to evaluate (4.33), where the right hand side is an expression depending on the complex structure coordinates $(x_{1},x_{2})$, for which we eventually take the limit to $P_{+}$, i.e. $x_{1}\to 0$ for fixed $x_{2}$. More generally, we consider the limit $x_{1}\to 0$ along a generic path $x_{2}(x_{1})=c+O(x_{1})$ with the constant $c\neq 0$.
Similar to the computation for the Quintic, we have to keep track of the leading singularities and the combinations of logarithmic factors with linear terms in the $x_{1}$. Doing this, we realize that we can express the derivatives of the Kähler moduli as
$\begin{split}{\partial t_{i}\over\partial x_{1}}=(\log x_{1})\,{\cal G}^{(i1)}(x)\,,\qquad\quad{\partial t_{i}\over\partial x_{2}}=(x_{1}\log x_{1})\,{\cal G}^{(i2)}(x)\ \end{split}$ (4.36)
with non-singular functions ${\cal G}^{(ij)}(x)$ having a general expansion
$\begin{split}{\cal G}^{(i,j)}(x_{1},x_{2})=\sum_{m,n,p=0}^{\infty}{(x_{1}\log x_{1})^{m}\over\log^{p}(x_{1})}\left({\log x_{2}\over\log x_{1}}\right)^{n}a_{m,n,p}^{(ij)}\,,\end{split}$ (4.37)
where the $a_{m,n,p}^{(ij)}$ are polynomials in $(x_{1},x_{2})$ having a non-vanishing constant term. In (4.36) most of the $a_{m,n,p}^{(ij)}$ are actually vanishing but in the next step we have to invert such expressions and then, like for the Quintic, we generate the full infinite series. Note that in the limit $x_{1}\to 0$ and fixed $x_{2}$ all terms in (4.37) are controlled, i.e.
$\begin{split}(x_{1}\log(x_{1}))^{m}\to 0\,,\qquad\log^{-p}(x_{1})\to 0\,,\qquad\left({\log x_{2}\over\log x_{1}}\right)^{n}\to 0\,.\end{split}$ (4.38)
Now, inverting the Jacobian and using (4.34) we insert everything into (4.33), to realize that we can write the final results as a sum of four terms
$\begin{split}\partial_{t_{i}}\partial_{t_{j}}\partial_{t_{k}}\mathcal{F}_{0}|_{\rm weak}=&\frac{1}{x_{1}\log^{3}(x_{1})}\,{\cal G}_{1}^{(ijk)}(x)+\frac{1}{x_{1}\log^{3}(x_{1})}\,{\cal G}_{2}^{(ijk)}(x)\\[5.69046pt] &+\frac{1}{\log^{3}(x_{1})}\,{\cal G}_{3}^{(ijk)}(x)+\frac{1}{x_{1}\log^{3}(x_{1})}\,{\cal G}_{4}^{(ijk)}(x)\,.\end{split}$ (4.39)
Here the ${\cal G}^{(ijk)}_{n}(x)$ are again functions of the type (4.37). The first, the second and the last term show the familiar conifold singularity. The third term directly vanishes in the $x_{1}\to 0$ limit. As for the Quintic, it is now evident that independent of the “angle” $x_{2}$, in the limit $x_{1}\to 0$ there are no constant terms arising from the expression (4.39). There are only divergent terms, which we subtract and terms that go to zero. Hence, applying the regularization (4.29) one correctly obtains the four TINs
$\begin{split}Y_{t_{1}t_{1}t_{1}}^{(0)}=9\,,\qquad Y_{t_{1}t_{1}t_{2}}^{(0)}=3\,,\qquad Y_{t_{1}t_{2}t_{2}}^{(0)}=1\,,\qquad Y_{t_{2}t_{2}t_{2}}^{(0)}=0\,.\end{split}$ (4.40)
One might be worried about taking the limit towards the degeneration point $P_{+}$ along a different family of paths and getting a different result. This can indeed be the case and we will come back to this issue in due course. Moreover, we have obtained the desired results by taking the limit towards the intersection point $P_{+}$, which after resolving the moduli space is just one of the potential degeneration points (see [33]). For $P_{-}$ we expect that the computation goes through completely analogous to $P_{+}$ However, for handling the triple intersection point $P_{3}=D_{1}\cap D_{2}\cap D_{0}$ one has to perform the resolution shown already in figure 1. How this is done locally is explained in [46]. Since we have already identified a working intersection point, we are not doing these tedious computations for this CY but will present an example for $\mathbb{P}_{1,1,2,2,6}[12]$.
4.3 Emergence of TINs on $\mathbb{P}_{1,1,2,2,6}[12]$
As a second example with two Kähler moduli let us consider the CY manifold $\mathbb{P}_{1,1,2,2,6}[12]$ with Hodge numbers $(h_{11},h_{21})=(2,128)$. The GLSM data specifying the resolved threefold are given in table 4 (see [30]).
$z_{1}$ $z_{2}$ $z_{3}$ $z_{4}$ $z_{5}$ $z_{6}$ $p$
$l^{(1)}$ 0 0 1 1 3 1 -6
$l^{(2)}$ 1 1 0 0 0 -2 0
Table 4: Data specifying $K3$-fibered CY
The threefold is a $K3$-fibration with base $\mathbb{P}^{1}$, where $z_{1}=0$ defines a divisor $L$ of topology $K3$ and $z_{3}=0$ a second divisor $H$, so that for this basis one gets the TINs [45]
$\begin{split}H^{3}=4\,,\quad H^{2}\cdot L=2\,,\quad H\cdot L^{2}=0\,,\quad L^{3}=0\,.\end{split}$ (4.41)
Expanding the Kähler form as $J=t_{1}\,\omega_{H}+t_{2}\,\omega_{L}$ one finds that $t_{2}$ measures the size of the $\mathbb{P}^{1}$ base and that the Kähler cone is $t_{1},t_{2}>0$.
As a $K3$-fibration there exists a heterotic dual model so that the corresponding emergent string limit is approached for large radius of the base. Then the two complex structure moduli of the mirror dual CY are parametrized by the deformations of the hypersurface constraint
$\begin{split}p=z_{1}^{12}+z_{2}^{12}+z_{3}^{6}+z_{4}^{6}+z_{5}^{2}-12\psi\,z_{1}z_{2}z_{3}z_{4}z_{5}-2\phi\,z_{1}^{6}z_{2}^{6}\,,\end{split}$ (4.42)
where it is customary to introduce the combinations
$\begin{split}\overline{x}=-\frac{1}{2^{5}3^{3}}{\phi\over\psi^{6}}\,,\qquad\overline{y}={1\over\phi^{2}}\,.\end{split}$ (4.43)
In terms of these, the manifold features two degeneration loci at
$\begin{split}\Delta_{1}&=(1-\overline{x})^{2}-\overline{x}^{2}\overline{y}=0\,,\\ \Delta_{2}&=1-\overline{y}=0\end{split}$ (4.44)
where we notice that only the first one is a conifold singularity. The LCS point is at $\overline{x}=\overline{y}=0$ and one can solve the corresponding Picard-Fuchs equations in the vicinity of this point, determine an integral symplectic basis of periods $\Pi^{\rm LCS}_{I}$, $I=1,\ldots,6$, such that the mirror map is
$\begin{split}t_{1}={\Pi^{\rm LCS}_{2}\over\Pi^{\rm LCS}_{1}}={1\over 2\pi i}\log(\overline{x})+\ldots\,,\qquad t_{2}={\Pi^{\rm LCS}_{3}\over\Pi^{\rm LCS}_{1}}={1\over 2\pi i}\log(\overline{y})+\ldots\,.\end{split}$ (4.45)
Inverting the mirror map, the prepotential enjoys the familiar instanton expansion
$\begin{split}{\cal F}_{0}|_{\rm weak}={2\over 3}t_{1}^{3}+t_{2}t_{1}^{2}+{63i\zeta(3)\over 4\pi^{3}}+{\rm instantons}\,.\end{split}$ (4.46)
The exact B-side Yukawas were provided in [45, 33], which in terms of $(\bar{x},\bar{y})$ read [30]
$\begin{split}\kappa_{\bar{x}\bar{x}\bar{x}}&=\frac{2i}{\pi^{3}}\frac{1}{4\bar{x}^{3}((1-\bar{x})^{2}-\bar{x}^{2}\bar{y})}\,,\quad\phantom{aaaaaa}\kappa_{\bar{x}\bar{x}\bar{y}}=\frac{i}{2\pi^{3}}\frac{1-\bar{x}}{2\bar{x}^{2}\bar{y}((1-\bar{x})^{2}-\bar{x}^{2}\bar{y})}\,,\\ \kappa_{\bar{x}\bar{y}\bar{y}}&=\frac{i}{8\pi^{3}}\frac{2\bar{x}-1}{\bar{x}\bar{y}(1-\bar{y})((1-\bar{x})^{2}-\bar{x}^{2}\bar{y})}\,,\quad\kappa_{\bar{y}\bar{y}\bar{y}}=\frac{i}{32\pi^{3}}\frac{2(1-\bar{x}+\bar{y}-3\bar{x}\bar{y})}{\bar{y}^{2}(1-\bar{y})^{2}((1-\bar{x})^{2}-\bar{x}^{2}\bar{y})}\,.\end{split}$ (4.47)
4.3.1 Resolution of moduli space
For the following analysis we need to collect some more information about the resolution of the complex structure moduli space. For the emerging string limit, we are interested in the tangential intersection point between the conifold locus $D_{1}$ and the LCS locus $D_{\infty}$. As shown in figure 3, its resolution involves two exceptional divisors with now pairwise normal crossings. The local coordinates $(x_{1},x_{2})$ around the three intersections are [48]
$\begin{split}D_{\infty}\cap E_{1}:&\quad\bigg{(}{\overline{x}^{2}\overline{y}\over(1-\overline{x})^{2}},1-\overline{x}\bigg{)}\,,\\[2.84544pt] D_{1}\cap E_{1}:&\quad\bigg{(}{(1-\overline{x})^{2}-\overline{x}^{2}\overline{y}\over(1-\overline{x})^{2}},1-\overline{x}\bigg{)}\,,\\[2.84544pt] E_{2}\cap E_{1}:&\quad\bigg{(}{\overline{x}^{2}\overline{y}\over(1-\overline{x})},{(1-\overline{x})^{2}\over\overline{x}^{2}\overline{y}}\bigg{)}\,.\end{split}$ (4.48)
[…]
Figure 3: Schematic view of the intersections of degeneration loci $D_{1}=\{\Delta_{1}=0\}$, $D_{2}=\{\Delta_{2}=0\}$, $D_{0}$, $D_{\infty}$ and their resolution.
For the M-theory limit, the obvious intersection point $P_{1}$ of two degeneration loci $D_{1}$ and $D_{2}$ is at $(\overline{x},\overline{y})=(1/2,1)$, but, according to [45], there exists a second point $P_{2}$ at $(\overline{w}=\overline{x}^{-1},\overline{y})=(0,1)$, where they intersect also with a third orbifold locus $D_{0}$ located at $\psi=0$. To get a smooth moduli space with normal crossings of the degeneration loci, one needs to perform a resolution of this triple intersection. This is done via a blow-up, which introduces a divisor $E_{3}$ which has normal crossings with the three degeneration loci as shown on the right in figure 3. How this is done locally is explained in [46]. To describe the resolved complex structure moduli space around $P_{2}$, one needs three charts where each chart provides the local coordinates $(x_{1},x_{2})$ for the normal intersection of the divisor $E_{3}$ with one of the three degeneration loci $D_{1}$, $D_{2}$ and $D_{0}$
$\begin{split}P_{13}=D_{1}\cap E_{3}:&\quad\bigg{(}-{1\over\overline{x}}\big{(}(1-\overline{x})^{2}-\overline{x}^{2}\overline{y}\big{)},{1\over\overline{x}}\bigg{)}\,,\\ P_{23}=D_{2}\cap E_{3}:&\quad\bigg{(}\bar{x}(1-\overline{y}),\frac{1}{\overline{x}}\bigg{)}\,,\\ P_{03}=D_{0}\cap E_{3}:&\quad\bigg{(}\frac{1}{\overline{x}(1-\overline{y})},1-\overline{y}\bigg{)}\,.\end{split}$ (4.49)
Note that in all three charts, the point $P_{2}$ with $\overline{w}_{0}=1/\overline{x}_{0}=0$ and $\overline{y}_{0}=1$ is reached for $x_{2}=0$ and $x_{1}$ can be regarded as an angular variable.
4.3.2 Emergent heterotic string limit
For large $t_{2}$ this model has a weakly coupled heterotic dual on $K3\times T^{2}$, where the heterotic 4d complexified dilaton and the Kähler parameter of the $T^{2}$ factor are related to the two Kähler moduli of the type IIA CY as
$\begin{split}t_{2}=iS\,,\qquad t_{1}=iT\,.\end{split}$ (4.50)
Hence, $\overline{y}\to 0$ corresponds to the weak coupling limit on the heterotic side. In the heterotic prepotential
$\begin{split}{\cal F}_{0}^{\rm het}=i\bigg{(}-ST^{2}-\left({2\over 3}T^{3}+O(e^{-2\pi T})\right)+O(e^{-2\pi S})\bigg{)}\,,\end{split}$ (4.51)
the first term is at tree-level, the second one the one-loop correction and the last one a sum over heterotic $NS5$-brane instantons. Now, let us analyze how (part of) the prepotential is generated in strongly coupled type IIA limits.
First, we take the emergent string limit on the type IIA side, i.e. as in section 3.2 we co-scale the size of base like $t_{2}\to\lambda t_{2}$ and the type IIA string coupling as $g^{\rm IIA}_{s}\to\lambda^{1/2}g^{\rm IIA}_{s}$ so that the 4D Planck-scale remains constant. This is the same limit as for the STU model so that for the same reason not the full set of Yukawa-couplings is emerging from integrating out wrapped $D2$-$D0$ bound states at one loop. In fact, emergence in this case means that the TIN $\kappa_{t_{1}t_{1}t_{1}}=4$ is given by the regularization of the zero point Yukawa coupling
$\begin{split}Y_{t_{1}t_{1}t_{1}}^{(0)}=-\lim_{t_{1}\to t_{1,c}}\lim_{t_{2}\to\infty}\left[\Big{(}\partial^{3}_{t_{1}}{\cal F}_{0}\big{|}_{\rm weak}-\kappa_{t_{1}t_{1}t_{1}}\Big{)}-{\rm Div}\right]\,.\end{split}$ (4.52)
This is nothing else than the constant piece in the one-loop correction to the Yukawa coupling (4.51) on the dual heterotic side. The first limit $t_{2}\to\infty$ guarantees that before regularizing the zero point Yukawa coupling via the limit $\lim_{t_{1}\to t_{1,c}}$, all $D2$-branes wrapping the large base $\mathbb{P}^{1}$ decouple.
This means that on the mirror dual side, we need to take the limit to the point $(\overline{x},\overline{y})=(1,0)$, which is the intersection of the conifold locus and the LCS locus in $\overline{y}$. Therefore, to actually evaluate the expression (4.52) we need to know $\partial^{3}_{t_{1}}{\cal F}_{0}\big{|}_{\rm weak}$ in a vicinity of this point in complex structure moduli space. Since, the two loci intersect only tangentially one needs to perform the resolution described in section 4.3.1, which introduces three local charts with coordinates (4.48).
Let us consider the chart around $D_{\infty}\cap E_{1}$. Again, one first needs to solve the Picard-Fuchs equations in local coordinates and then to determine an integral symplectic basis which is required to be the continuation of the integral symplectic basis around the LCS point $(\overline{x},\overline{y})=(0,0)$. The resulting periods take the form
$\begin{split}X_{\rm 0}/\omega_{0}&=1\,\\ X_{\rm 1}/\omega_{0}&=P_{1}(x_{1},x_{2})\,\\ X_{\rm 2}/\omega_{0}&={1\over 2\pi i}\log(x_{1}x_{2}^{2})+{F_{\rm 2}/\omega_{0}\over 2\pi i}\log(x_{2})+P_{2}(x_{1},x_{2})\,\\ F_{\rm 2}/\omega_{0}&=\sqrt{x_{2}}P_{3}(x_{1},x_{2})\,\\ F_{\rm 1}/\omega_{0}&=-{X_{\rm 1}/\omega_{0}\over\pi i}\log(x_{1}x_{2}^{2})+P_{4}(x_{1},x_{2})\,\\ F_{\rm 0}/\omega_{0}&=-\frac{1}{\pi i}\log(x_{1}x_{2}^{2})-\frac{F_{2}/\omega_{0}}{2\pi i}\log(x_{1}x_{2}^{2})+P_{5}(x_{1},x_{2})\,,\end{split}$ (4.53)
where each $P_{i}(x_{1},x_{2})$ is a power series in $(x_{1},\sqrt{x_{2}})$. Indeed the periods $X_{2},F_{2}$ feature the expected conifold behavior. In the vicinity of $(x_{1},x_{2})=(0,0)$, the leading order behavior of the Kähler moduli is
$\begin{split}t_{1}&={X_{1}\over\omega_{0}}=i+c_{0}\sqrt{x_{2}}+\ldots\,,\\ t_{2}&={X_{2}\over\omega_{0}}={1\over 2\pi i}\log(x_{1}x_{2}^{2})+d_{0}+d_{1}\sqrt{x_{2}}\log(x_{2})+\ldots\,,\end{split}$ (4.54)
with $c_{i}$ and $d_{i}$ series in $x_{1}$ and $\sqrt{x_{2}}$. Next one can determine the Jacobian and insert its inverse into the expression for the third derivative of the prepotential. Eventually, in the weak heterotic string coupling limit $x_{1}\rightarrow 0$ this yields an expansion
$\begin{split}\partial_{t_{1}}^{3}{\cal F}_{0}={\,a_{-1}\over\sqrt{x_{2}}}+a_{0}\log\left({x_{2}\over\Lambda_{0}}\right)+\sum_{n=1}^{\infty}\,a_{n}\log\left({x_{2}\over\Lambda_{n}}\right)x_{2}^{n/2}\,,\end{split}$ (4.55)
where the $a_{n}$ and $\Lambda_{n}$ are real with $a_{-1}\approx 4.826$, $a_{0}\approx 1.910$ and $\Lambda_{0}\approx 0.320$. The first term diverges for the conifold limit $x_{2}\to 0$ and scales as
$\begin{split}Y_{t_{1}t_{1}t_{1}}^{(0)}\sim{1\over(t_{1}-i)}+\ldots\,.\end{split}$ (4.56)
In appendix A, we will relate this to the asymptotic behavior of the GV invariants. All the terms on the right hand side of (4.55) go to zero, while the middle term also diverges though also containing a constant term that is however ambiguous [^6] The constant piece in $\log(\epsilon A)$ depends on the scale, i.e. it can be changed by rescaling $\epsilon\to\lambda\epsilon$ and then taking $\epsilon\to 0$ and minimally subtract the divergences.. A very similar situation was also encountered in the regularization of the Schwinger integral for the $R^{4}$-term in 8D and the one-loop topological free energy in 4D, where such an ambiguous behavior was correlated with the appearance of conformal symmetry [17]. Hence, we consider such ambiguous constants contained in $\log$-factors not as counter examples but as a sign that we are on the right track and might just not yet perform the computation in the best suitable chart. We indicate this by writing $Y_{t_{1}t_{1}t_{1}}^{(0)}=4^{*}$.
To substantiate this idea we next repeat the above analysis in the chart around $D_{1}\cap E_{1}$. The periods obey very similar relations and are given by
$\begin{split}X_{\rm 0}/\omega_{0}&=1\,\\ X_{\rm 1}/\omega_{0}&=R_{1}(x_{1},x_{2})\,\\ X_{\rm 2}/\omega_{0}&=\frac{1}{2\pi i}\log(x_{2}^{2})+\frac{F_{2}/\omega_{0}}{2\pi i}\log(x_{2})+R_{2}(x_{1},x_{2})\,\\ F_{2}/\omega_{0}&=\sqrt{x_{2}}\,R_{3}(x_{1},x_{2})\,\\ F_{1}/\omega_{0}&=-\frac{X_{1}/\omega_{0}}{\pi i}\log(x_{2}^{2})+R_{4}(x_{1},x_{2})\,\\ F_{0}/\omega_{0}&=-\frac{1}{\pi i}\log(x_{2}^{2})-\frac{F_{2}/\omega_{0}}{\pi i}\log(x_{2})+R_{5}(x_{1},x_{2})\,.\\ \end{split}$ (4.57)
Here each series $R_{i}(x_{1},x_{2})$ has the expansion
$\begin{split}R_{i}(x_{1},x_{2})=\sum_{n=0}^{\infty}(x_{1}\sqrt{x_{2}}\log(x_{1}))^{n}r^{(i)}_{n}(x_{1},x_{2})\,.\end{split}$ (4.58)
The $r^{(i)}_{n}$ are polynomials in $x_{1}$ and $\sqrt{x_{2}}$ and start with a constant. In the vicinity of $(x_{1},x_{2})=(0,0)$, the leading terms in the Kähler moduli are
$\begin{split}t_{1}&=\frac{X_{1}}{\omega_{0}}=i+c_{0}\sqrt{x_{2}}\,+c_{1}x_{1}\sqrt{x_{2}}\,\log(x_{1})+\ldots\,,\\ t_{2}&=\frac{X_{2}}{\omega_{0}}=\frac{1}{2\pi i}\log(x_{2}^{2})+d_{0}+d_{1}\sqrt{x_{2}}\,\log(x_{2})+\ldots\,.\end{split}$ (4.59)
Also here $c_{i}$ and $d_{i}$ denote series in $x_{1},\sqrt{x_{2}}$. Once again we want to evaluate $\partial_{t_{1}}^{3}{\cal F}_{0}$ in the heterotic string weak coupling limit $\bar{y}\rightarrow 0$. In the present chart the coordinates satisfy the relation $\bar{x}^{2}\bar{y}=x_{2}^{2}(1-x_{1})$. In the limit $x_{1}\rightarrow 1$ we reach the boundary of this chart, where we cannot trust the above periods any more. This suggests that we should rather take $x_{2}\rightarrow 0$ (keeping $x_{1}$ small and constant), which according to (4.59) already ensures that $t_{2}\rightarrow i\infty$ and $t_{1}\rightarrow t_{1}^{(0)}=i$. The expansion of $\partial_{t_{1}}^{3}{\cal F}_{0}$ for small $x_{1},x_{2}$ yields
$\begin{split}\partial_{t_{1}}^{3}{\cal F}_{0}=\frac{f_{0}(x_{1})}{x_{1}\sqrt{x_{2}}\log^{3}(x_{1})}+\frac{f_{1}(x_{1})+f_{2}(x_{1})\log(x_{2})}{x_{1}\log^{3}(x_{1})}+\ldots\,,\end{split}$ (4.60)
where $f_{i}(x_{1})$ are functions of $x_{1}$ and ellipses denote terms that vanish for $x_{2}\rightarrow 0$. We find the same leading singularity in the variable $x_{2}$ as in (4.55). Moreover, there is a series of $\log(x_{2})$-divergent terms that are paired with constants similar to the middle term in (4.55). However, unlike in the previous case these constants all depend on $x_{1}$, but treating $x_{1}$ as a constant we essentially get the same result as for the chart $D_{\infty}\cap E_{1}$.
However, the absence of a bare, i.e. not $x_{1}$ dependent, constant in (4.60) makes it possible to define a certain family of paths towards the singularity where $x_{1}$ scales as well. If one for example chooses $x_{1}=z\ \sqrt{x_{2}}$ (with $z$ constant), no constant term remains after minimally subtracting the divergences and taking the limit $x_{2}\rightarrow 0$. Hence, the absence of a bare constant allowed us to find a family of paths that indeed yields the desired value of the regularized Yukawa coupling $Y_{t_{1}t_{1}t_{1}}^{(0)}=4$.
4.3.3 M-theory limit I
So far all the limits we were considering for the two CYs with two Kähler moduli were approaching conifold singularities. For these the proposed regularization condition was satisfied. In this respect the M-theory limit for $\mathbb{P}_{1,1,2,2,6}[12]$ is interesting, as it will involve taking the limit to the intersection $\Delta_{1}\cap\Delta_{2}$ of the two degeneration limits, where $\Delta_{2}$ is not a conifold locus, but a strong coupling locus. The obvious intersection point $P_{1}$ is at $(\overline{x},\overline{y})=(1/2,1)$, but, as we have seen, there exists a second point $P_{2}$ at $(\overline{w}=\overline{x}^{-1},\overline{y})=(0,1)$, where they intersect also with a third orbifold locus at $\psi=0$. The question is whether all four TINs do indeed arise from the regularization of the corresponding zero-point Yukawa couplings $Y^{(0)}_{ijk}$.
To evaluate the limit towards $P_{1}$, we need to determine the integral symplectic basis of periods around the intersection point $(\overline{x},\overline{y})=(1/2,1)$. For that purpose it is useful to introduce the local coordinates
$\begin{split}x_{1}=1-2\overline{x}\,,\qquad\quad x_{2}={1-\overline{y}\over 1-2\overline{x}}\,,\end{split}$ (4.61)
where like for the coordinates of $\mathbb{P}_{1,1,1,6,9}[18]$ around $P_{+}$, $x_{2}$ is an angular variable and the singularity is reached by taking the limit $x_{1}\rightarrow 0$ and $x_{2}=c\neq 0$ kept constant. Then the two degeneration loci are at
$\begin{split}\Delta_{1}=x_{1}\big{(}1+O(x_{i})\big{)}\,,\qquad\quad\Delta_{2}=x_{1}x_{2}\,.\end{split}$ (4.62)
The periods take the form
$\begin{split}X_{\rm 0}/\omega_{0}&=1\,\\[2.84544pt] X_{\rm 1}/\omega_{0}&=-\frac{1}{2}X_{2}+S_{1}(x)\,\\[2.84544pt] X_{\rm 2}/\omega_{0}&=\sqrt{x_{1}x_{2}}\,S_{2}(x)\,\\[2.84544pt] F_{\rm 2}/\omega_{0}&=\frac{1}{2}F_{1}-\frac{1}{2\pi i}X_{2}\log(x_{2})+\sqrt{x_{1}x_{2}}\,S_{3}(x)\,\\[2.84544pt] F_{\rm 1}/\omega_{0}&=S_{4}(x)\\[2.84544pt] F_{\rm 0}/\omega_{0}&=x_{1}S_{5}(x)\,,\end{split}$ (4.63)
where $(x)$ denotes dependence of $(x_{1},x_{2})$ and each of the series $S_{i}$ has the general expansion
$\begin{split}S_{i}(x)=\sum_{n=0}^{\infty}(x_{1}\log(x_{1}))^{n}s_{n}^{(i)}(x)\end{split}$ (4.64)
and the $s_{n}^{(i)}$ are polynomials in $(x_{1},x_{2})$ starting with a constant. In the present case the two Kähler moduli share common terms and enjoy the expansion
$\begin{split}t_{1}=S_{1}(x_{1},x_{2})-\frac{\sqrt{x_{1}x_{2}}}{2}\,S_{2}(x_{1},x_{2})\,,\qquad\quad t_{2}=\sqrt{x_{1}x_{2}}\,S_{2}(x_{1},x_{2})\,.\end{split}$ (4.65)
In particular, for both $S_{1}$ and $S_{2}$ each $x_{2}^{n}$ term is again always accompanied by at least one factor of $x_{1}$. In the limit $x_{1}\rightarrow 0$ one obtains
$\begin{split}t^{(0)}_{1}\approx 1.334i\,,\qquad\quad t^{(0)}_{2}=0\,.\end{split}$ (4.66)
Note that in contrast to the $\mathbb{P}_{1,1,1,6,9}[18]$ case, here we see only the typical conifold behavior in $x_{1}$. The other degeneration at $\Delta_{1}$ only leads to power-law behavior. Close to $x_{1}=0$ the two degeneration loci are mapped to $t_{2}=-2(t_{1}-t_{1}^{(0)})$ and $t_{2}=0$ respectively. This means that close to $x_{1}=0$, the mirror dual of the LCS phase is the cone
$\begin{split}C_{\rm LCS}=\Big{\{}{\rm Im}(t_{2})>0,{\rm Im}(t_{1})>t_{1}^{(0)}-{1\over 2}{\rm Im}(t_{2})\Big{\}}.\end{split}$ (4.67)
Note that compared to the $\mathbb{P}_{1,1,1,6,9}[18]$ there is an essential difference when approaching the singular point in the Kähler moduli space. From (4.65) one can directly infer that for $x_{2}$ constant and small $x_{1}$ one gets
$\begin{split}{\Delta t_{2}\over\Delta t_{1}}=-2+O(x_{1})\end{split}$ (4.68)
so that for any fixed value of $x_{2}$ one approaches the singularity in a tangential manner. This also evident from the parametric plot in figure 4, where we observe that the shaded region is inside the cone $C_{\rm LCS}$.
Figure 4: Parametric plot of $({\rm Im}(t_{1}-t_{1}^{(0)}),{\rm Im}(t_{2}))$ for varying $(x_{1},x_{2})$ with ${\rm Im}(x_{i})=0$ and $0\leq{\rm Re}(x_{1})\leq 0.1$, $0.1\leq{\rm Re}(x_{2})\leq 1$.
Now we collect the remaining ingredients. Transforming the Yukawa couplings (4.47) to the new coordinates yields
$\begin{split}\kappa_{x_{1}x_{1}x_{1}}&={x_{2}\over x^{2}_{1}}P_{1}(x_{1},x_{2})\,,\qquad\kappa_{x_{1}x_{1}x_{2}}=\frac{1}{x_{1}}P_{2}(x_{1},x_{2})\,,\\ \kappa_{x_{1}x_{2}x_{2}}&=\frac{1}{x_{2}}P_{3}(x_{1},x_{2})\,,\qquad\;\kappa_{x_{2}x_{2}x_{2}}={x_{1}\over x_{2}^{2}}P_{4}(x_{1},x_{2})\,.\end{split}$ (4.69)
Once again, the $P_{i}$ are polynomials in $(x_{1},x_{2})$ with non-vanishing constant term.
Next, we need to determine the leading behavior of the partial derivatives $\partial x_{\alpha}/\partial t_{j}$. The derivatives of the mirror map scale as
$\begin{split}\frac{\partial t_{1}}{\partial x_{1}}&=\log(x_{1})\,\mathcal{G}_{1}(x)+\sqrt{x_{2}\over x_{1}}\,\mathcal{H}_{1}(x)\,,\qquad\frac{\partial t_{1}}{\partial x_{2}}=\sqrt{x_{1}\over x_{2}}\,\mathcal{G}_{2}(x)\,,\\[2.84544pt] \frac{\partial t_{2}}{\partial x_{1}}&=\sqrt{x_{2}\over x_{1}}\,\mathcal{G}_{3}(x)\,,\qquad\phantom{aaaaaaaaaaaaia}\frac{\partial t_{2}}{\partial x_{2}}=\sqrt{x_{1}\over x_{2}}\,\mathcal{G}_{4}(x)\,\end{split}$ (4.70)
with $\mathcal{G}_{i}$ and $\mathcal{H}$ all regular and finite in the limit $x_{1}\rightarrow 0$ with $x_{2}$ constant. The inverse of the Jacobian matrix again yields the derivatives $\partial x_{\alpha}/\partial t_{j}$ as functions of $(x_{1},x_{2})$. Quite remarkably, the leading-order term of the Jacobian determinant is not a constant, as one would naively expect from (4.70). The relations (4.65) among the Kähler moduli lead to cancellations, so that the determinant scales as
$\begin{split}\det J_{\rm mir}=\sqrt{x_{1}}\log(x_{1})\,\mathcal{G}_{\rm det}(x)\,,\end{split}$ (4.71)
where $\mathcal{G}_{\rm det}$ is also regular in the limit we take.
Finally, we combine all the ingredients to obtain the Yukawa couplings on the A-side. Evaluating (4.29) one obtains that for two out of the four regularized zero point Yukawa couplings one indeed finds no constant in the limit of $\partial_{t_{i}}\partial_{t_{j}}\partial_{t_{k}}\mathcal{F}_{0}|_{\rm weak}$ so that
$\begin{split}Y_{t_{1}t_{1}t_{1}}^{(0)}=4\,,\qquad\quad Y_{t_{1}t_{1}t_{2}}^{(0)}=2\,.\end{split}$ (4.72)
However, for the other two, actually vanishing TINs the regularization yields finite constant contributions so that
$\begin{split}Y_{t_{1}t_{2}t_{2}}^{(0)}\approx 0.24\,,\qquad\quad Y_{t_{2}t_{2}t_{2}}^{(0)}\approx 0.36\,.\end{split}$ (4.73)
Hence, the proposed regularization method is only successful in half of the cases. For the two correctly regularized zero-point Yukawa couplings (4.72), the generic scaling of their maximal divergence for $x_{1}\to 0$ is
$\begin{split}Y^{(0)}\sim{1\over x_{1}\log^{3}(x_{1})}+\ldots\sim{1\over\big{(}t_{1}-t_{1}^{(0)}+{1\over 2}t_{2}\big{)}\log^{2}\big{(}t_{1}-t_{1}^{(0)}+{1\over 2}t_{2}\big{)}}+\ldots\,.\end{split}$ (4.74)
In appendix A we investigate whether this behavior can be directly related to the asymptotic growth of the Gopakumar-Vafa invariants.
Looking at figure 4, one might suspect that the mismatch of (4.73) is related to the non-generic, tangential limit that we are taking. Then we should be able to improve the results by finding more generic paths towards the degeneration locus. In the concrete case, such paths can be found by balancing the two contributions for $\partial t_{1}/\partial x_{1}$ in (4.70). Approaching the singularity along paths with
$\begin{split}\sqrt{x_{1}x_{2}}=-zx_{1}\log(x_{1})\,,\end{split}$ (4.75)
we find that indeed the family of paths is now more generic but the four resulting regularized zero point Yukawa couplings
$\begin{split}Y_{t_{1}t_{1}t_{1}}^{(0)}=4\,,\qquad Y_{t_{1}t_{1}t_{2}}^{(0)}=2^{\star}\,,\qquad Y_{t_{1}t_{2}t_{2}}^{(0)}=0^{\star}\,,\qquad Y_{t_{2}t_{2}t_{2}}^{(0)}\approx 0.36+{0.15\over z}\end{split}$ (4.76)
are still not all correct. Here the star again indicates that this is only up to conformal scaling. As in the previous section 4.3.2, the presence of bare constants in (4.73) only allows us to make some of them ambiguous but not to get rid of them completely.
4.3.4 M-theory limit II
Next we consider taking the limit towards the singular point $P_{2}$ located at $(\bar{x}=\infty\,,\bar{y}=1)$, where the two degeneration loci $\Delta_{1}$ and $\Delta_{2}$ intersect with the singular curve $D_{0}$ that arises from an orbifold action on the moduli space coordinates. As discussed in section 4.3.1, several blow-ups are required to achieve normal crossings between those divisors so that, after the resolution, there are three points of intersection $\mathcal{P}\in\{P_{13},P_{23},P_{03}\}$.
We will now analyze the point $P_{03}$, which is the intersection of $D_{0}$ with the exceptional divisor $E_{3}$ separating the curves parametrized by $\Delta_{1}$ and $\Delta_{2}$. To obtain the integral symplectic basis we solve the Picard-Fuchs equations centered around each of the three above points and match the periods in suitable transition regions step by step. Around $(x_{1},x_{2})=(0,0)$, the periods take the form
$\begin{split}X_{0}&=(x_{1}x_{2})^{1/6}A_{0}+\sqrt{x_{1}x_{2}}\log(x_{2})B_{0}\\ X_{1}&=(x_{1}x_{2})^{1/6}A_{1}+\sqrt{x_{1}x_{2}}\log(x_{2})B_{1}\\ X_{2}&=x_{1}^{1/6}x_{2}^{1/2}C_{2}\\ F_{0}&=(x_{1}x_{2})^{1/6}\hat{A}_{0}+\sqrt{x_{1}x_{2}}\log(x_{2})\hat{B}_{0}\\ F_{1}&=(x_{1}x_{2})^{1/6}\hat{A}_{1}+\sqrt{x_{1}x_{2}}\log(x_{2})\hat{B}_{1}\\ F_{2}&=(x_{1}x_{2})^{1/6}\hat{A}_{2}+\sqrt{x_{1}x_{2}}\log(x_{2})\hat{B}_{2}\,,\end{split}$ (4.77)
where $A_{i},B_{i},\hat{A}_{i},\hat{B}_{i}$ denote polynomials in $(x_{1}^{1/3},x_{2}^{1/3})$ and each $x_{1}^{n/3}$ is paired at least with a factor of $x_{2}^{1/3}$. $C_{2}$ is a polynomial in $(x_{1}^{1/3},x_{2})$ and contains bare $x_{1}$-terms. The Kähler moduli are given by the expansions
$\begin{split}t_{1}&=\sum_{n=0}^{\infty}\big{(}x_{1}^{1/3}x_{2}^{1/3}\log(x_{2})\big{)}^{n}a_{n}=t_{1}^{(0)}+x_{2}^{1/3}c_{0}+x_{1}^{1/3}x_{2}^{1/3}\log(x_{2})\,c_{1}+\ldots\,,\\[2.84544pt] t_{2}&=x_{2}^{1/3}\sum_{n=0}^{\infty}\big{(}x_{1}^{1/3}x_{2}^{1/3}\log(x_{2})\big{)}^{n}b_{n}=t_{2}^{(0)}+x_{2}^{1/3}d_{0}+x_{1}^{1/3}x_{2}^{2/3}\log(x_{2})\,d_{1}+\ldots\,.\end{split}$ (4.78)
Here $a_{n},b_{n},c_{n},d_{n}$ are also series in $(x_{1}^{1/3},x_{2}^{1/3})$. In the limit $x_{2}\rightarrow 0$ we obtain [^7] Of course, for $t_{1}^{(0)}$ we only obtain numerical values that are suspiciously close (up to order $O(10^{-5})$) to the presented ones.
$\begin{split}t_{1}^{(0)}=-{1\over 2}+{\sqrt{3}\over 2}i\,,\qquad\qquad t_{2}^{(0)}=0\,.\end{split}$ (4.79)
This limiting value is common for all three points from $\mathcal{P}$ when $x_{2}$ is taken to zero, so that indeed we have three pathches around a common intersection point $P_{0}$. This is also shown in figure 5.
Figure 5: Parametric plot of $({\rm Im}(t_{1}-t_{1}^{(0)}),{\rm Im}(t_{2}))$, evaluated in the patches around $\{P_{23},P_{13},P_{03}\}$. We vary the respective coordinates $(x_{1},x_{2})$ with ${\rm Im}(x_{1})=0$, ${\rm Im}(x_{2})=0$. For $P_{23}$ we have $10.1<{\rm Re}(x_{1})<$ 1 and $0.10<{\rm Re}(x_{2})<$ 0.1, for $P_{13}$ it’s $10.01<{\rm Re}(x_{1})<$ 1 and $0.050<{\rm Re}(x_{2})<$ 0.05 and for $P_{03}$ we choose $0.50.1<{\rm Re}(x_{1})<$ 0.5 and $0.10<{\rm Re}(x_{2})<$ 0.1.
To determine the Yukawa couplings we proceed as for the previous cases. Transforming the B-side Yukawas to the coordinates of the patch around $P_{03}$ yields
$\begin{split}\kappa_{x_{1}x_{1}x_{1}}&={x_{2}\over x_{1}}P_{1}(x_{1},x_{2})\,,\qquad\kappa_{x_{1}x_{1}x_{2}}=\frac{1}{x_{1}}P_{2}(x_{1},x_{2})\,,\\ \kappa_{x_{1}x_{2}x_{2}}&=\frac{1}{x_{2}}P_{3}(x_{1},x_{2})\,,\qquad\;\kappa_{x_{2}x_{2}x_{2}}={x_{1}\over x_{2}^{2}}P_{4}(x_{1},x_{2})\,,\end{split}$ (4.80)
where each $P_{i}$ is a polynomial in $(x_{1},x_{2})$. We compute the partial derivatives $\partial x_{\alpha}/\partial t_{j}$ again by using the derivatives of the mirror map, which are given by
$\begin{split}\frac{\partial t_{1}}{\partial x_{1}}&=\frac{x_{2}^{1/3}}{x_{1}^{2/3}}\log(x_{2})\,\mathcal{G}_{1}(x)\,,\\ \frac{\partial t_{1}}{\partial x_{2}}&=\frac{1}{x_{2}^{2/3}}\mathcal{G}_{2}(x)\ +\frac{x_{1}^{1/3}}{x_{2}^{2/3}}\log(x_{2})\,\mathcal{H}_{2}(x)+\frac{x_{1}^{2/3}}{x_{2}^{1/3}}\log^{2}(x_{2})\,\mathcal{I}_{2}(x)+x_{1}\log^{3}(x_{2})\mathcal{J}_{2}(x)\,,\\ \frac{\partial t_{2}}{\partial x_{1}}&=\frac{x_{2}^{1/3}}{x_{1}^{2/3}}\,\mathcal{G}_{3}(x)\,,\\ \frac{\partial t_{2}}{\partial x_{2}}&=\frac{1}{x_{2}^{2/3}}\,\mathcal{G}_{4}(x)+\frac{x_{1}^{1/3}}{x_{2}^{1/3}}\log(x_{2})\,\mathcal{H}_{4}(x)+x_{1}^{2/3}\log^{2}(x_{2})\,\mathcal{I}_{4}(x)\,,\\ \end{split}$ (4.81)
where each of the $\mathcal{G}_{i},\mathcal{H}_{i},\mathcal{I}_{i},\mathcal{J}_{i}$ are regular and finite in the limit $x_{2}\rightarrow 0$. The leading divergence of the Jacobian determinant scales like
$\begin{split}\det J_{\rm mir}=\frac{\log(x_{2})}{x_{1}^{2/3}x_{2}^{1/3}}\,\mathcal{G}_{\rm det}\,,\end{split}$ (4.82)
with $\mathcal{G}_{\rm det}$ also finite and non-divergent as before.
Finally, we combine all these results to evaluate the A-side zero point Yukawa couplings. We observe that their divergence only appears from the scaling $X_{0}\sim(x_{1}x_{2})^{1/6}$ of the fundamental period. Following the natural family of paths with $x_{1}=c\neq 0$ and $x_{2}\rightarrow 0$, we find no constant for two out of four regularized zero point Yukawa couplings, namely
$\begin{split}Y_{t_{1}t_{1}t_{1}}^{(0)}=4\,,\qquad Y_{t_{1}t_{1}t_{2}}^{(0)}=2\,.\end{split}$ (4.83)
For the remaining two Yukawa couplings we get no bare constants, but a series of $x_{1}$-dependent terms that survive in the specified limit. Hence, we essentially arrive that the same result as for the M-theory limit I.
However, in contrast to the M-theory limit I, the absence of a bare constant and the experience from section 4.3.2 suggest that one can define another family of paths that works. One such family is $x_{1}=z\,x_{2}$ ($z$ constant), which ensures that both Kähler moduli in (4.78) scale as $x_{2}^{1/3}$ at leading order. In this case all regularized zero point Yukawa couplings carry no constant and really give the TINs
$\begin{split}Y_{t_{1}t_{1}t_{1}}^{(0)}=4\,,\qquad Y_{t_{1}t_{1}t_{2}}^{(0)}=2\,,\quad Y_{t_{1}t_{2}t_{2}}^{(0)}=0\,,\qquad Y_{t_{2}t_{2}t_{2}}^{(0)}=0\,.\end{split}$ (4.84)
The generic scaling of their divergence for $x_{2}\to 0$ is
$\begin{split}Y^{(0)}\sim{1\over x_{2}^{1/3}\log^{2}(x_{2})}\sim{1\over(t_{i}-t_{i}^{(0)})\log^{2}(t_{i}-t_{i}^{(0)})}\,,\end{split}$ (4.85)
at least for three out of four zero point Yukawa couplings.
The lesson we draw from this example is that the proposed regularization procedure is, maybe not surprisingly, sensitive to the actual co-dimension two degeneration point and the family of paths taken towards it. In contrast to the previous $\mathbb{P}_{1,1,1,6,9}[18]$ CY, in this example, we had to approach not the point $P_{1}$ but $P_{2}$ to successfully obtain all four triple intersection numbers via our proposed regularization method of the zero point Yukawa couplings. The art is to find a maximal codimension degeneration point so that all regularized zero point Yukawa couplings do not contain any bare constant. Then, it seems to be possible to identify a family of paths so that the proposal is satisfied and the so defined regularization procedure gives the wanted set of TINs. The intuitive picture behind this is that by approaching the complex codimension 2 degeneration locus, the limit has to be sensitive to the asymptotic growth of the GV invariants along all directions of the two-dimensional homology lattice. From the few examples studied so far, we could not yet identify a clear pattern that allows us to decide beforehand what the best suited patch around each degeneration point is.
5 Conclusions
The M-theoretic Emergence Proposal asserts that in the isotropic M-theory limit of type IIA, all interactions are generated via quantum effects. This means that also the type IIA string tree-level terms need to arise quantum mechanically. This has been explicitly demonstrated for higher derivative interactions, like $R^{4}$ and $F^{4}$ terms, in maximally respectively half-maximally supersymmetric theories, where they arose at one-loop level from a real Schwinger integral, after an appropriate regularization of the appearing UV divergences had been implemented.
In this paper, we were trying to fill a gap from [14], where a similar result was obtained for the type IIA string tree-level contribution to the $N=2$ prepotential in a 4D compactification of type IIA on a non-compact CY threefold, namely for the resolved conifold. There, the ultimate question remained, how the classical TINs could be obtained from a regularization of the GV-type one-loop Schwinger integral for compact CYs. The problem is that even after regularizing the Schwinger integral itself, eventually one has to carry out a highly diverging infinite sum over the GV invariants. Let us emphasize again that the tree-level kinetic terms for the type IIA Kähler moduli have their origin in the 10D Einstein-Hilbert term so that we are really dealing with the type of couplings mentioned in the original Emergence Proposal [1, 2, 3].
After concretizing the question and taking some lessons from a toy model, we approached this problem in a step by step procedure. It turned out to be helpful to not directly study the wanted M-theory limit but other infinite distance limits where only part of the cubic terms in the prepotential is expected to be generated at one-loop level. CY manifolds featuring an elliptic or $K3$- fibration structure were good candidates in this respect. Generalizing the findings to the isotropic M-theory limit, we managed to come up with an admittedly still speculative but mathematically concrete proposal for how the regularization of this infinite sums can be performed, namely
$\begin{split}Y^{(0)}_{t_{i}t_{j}t_{k}}=\frac{1}{2}\sum_{\beta\in H_{2}(X,\mathbb{Z})}^{\infty}\alpha_{0}^{\beta}\,\beta_{i}\beta_{j}\beta_{k}\,\Big{|}_{\rm reg}=-\lim_{t_{i}\to t_{i,0}}\left[\Big{(}\partial_{t_{i}}\partial_{t_{j}}\partial_{t_{k}}{\cal F}_{0}\big{|}_{\rm weak}-\kappa_{t_{i}t_{j}t_{k}}\Big{)}-{\rm Div}\right]\,.\end{split}$
We believe that it is satisfactory that the proposed regularization involves the TINs in an explicit way. Recall that the appearance of the weakly coupled type IIA prepotential is routed in the fact that, except for the triple intersection numbers to be subtracted explicitly, it implements the information about the infinitely many GV invariants [^8] A puzzling issue already mentioned in [18] is related to the Enriques CY, which is a free quotient $X=(K3\times T^{2})/\mathbb{Z}_{2}$ and known from [49] to have vanishing GV invariants $\alpha_{0}^{\beta}=0$. Clearly for this CY our regularization method is void and it remains unclear how the TINs can emerge from the now vanishing zero-point Yukawa couplings.. Effectively, we regularize the zero point Yukawa couplings via the world-sheet instanton contributions to the Yukawas and then take a small radius limit towards a codimension $h_{11}$ degeneration locus. For actually evaluating this expression we needed to determine the CY periods in a chart around the location $t_{i,0}$ of the singularity, i.e. information that we generated by solving the Picard-Fuchs equation and then continuously gluing these solutions to the periods in the well known large complex structure regime.
In this paper, we considered various degeneration loci of CYs with $h_{11}\leq 2$. First, we were studying the Quintic CY, which develops a conifold singularity at the degeneration locus (as do all 14 CYs with $h_{11}=1$ from the list (LABEL:listCYh11)). We were also investigating in some detail the degenerations of two CYs with $h_{11}=2$, namely the elliptic fibration $\mathbb{P}_{1,1,1,6,9}[18]$ and the $K3$ fibration $\mathbb{P}_{1,1,2,2,6}[12]$. The first one featured the intersection of two conifold loci and the second one the intersection of a conifold and a strong coupling locus. While for the (intersecting) conifold loci the regularization quite directly led to the wanted result $Y^{(0)}_{t_{i}t_{j}t_{k}}=\kappa_{t_{i}t_{j}t_{k}}$, for the $K3$ fibration the analysis turned out to be more involved. More specifically, the result of the above limit depended on the degeneration point and the local chart around it. We can summarize our findings in a refined regularization condition
For each CY there exist a codimension $h_{11}$ degeneration point $t_{i,0}$ and a local chart around it so that the $t_{i}\rightarrow t_{i,0}$ limit of $\partial_{t_{i}}\partial_{t_{j}}\partial_{t_{k}}\mathcal{F}_{0}|_{\rm weak}$ contains no bare constants. Then, we expect to find a family of paths so that the limit gives only divergent and vanishing terms. Hence, the so defined regularization of the zero-point Yukawa couplings yields $Y_{t_{i}t_{j}t_{k}}^{(0)}=\kappa_{t_{i}t_{j}t_{k}}$.
With only a few concrete examples studied in this paper, this should be considered as a working hypothesis for more detailed studies in the future, which should involve also CYs with more Kähler moduli and other degeneration loci.
Let us emphasize again that in this paper we were extending the successful regularization of UV-divergent real M-theoretic Schwinger integral from higher derivative terms with extended supersymmetry to the 4D prepotential in theories with $N=2$ supersymmetry. This is to be considered a pragmatic regularization method but lacks a truly finite microscopic description of the M-theoretic one-loop amplitudes. A proposal that goes in this direction has been put forward in [22, 23, 24] and it would be interesting to investigate any relation between these two complementary approaches. Nevertheless, we think that the investigation reported in this paper has revealed that the M-theoretic Emergence Proposal can be a powerful guide to reveal so far hidden structures in the relation between the asymptotic behavior of GV invariants and degenerations of CYs. One could entertain the picture that the moduli spaces of CY manifolds encode so much information about quantum gravity that their infinite distance degenerations are consistent with the emergent string conjecture and that their finite distance degenerations are consistent with the M-theoretic Emergence Proposal.
Acknowledgments:
We thank Niccolò Cribiori for collaboration on earlier stages of this work. Moreover, we are indebted to Rafael Álvarez-García for sharing his insights into period computations with us. We also acknowledge valuable discussions with Murad Alim and Antonia Paraskevopoulou. The work of R.B. and A.G. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
Appendix A Singularities and Gopakumar-Vafa invariants
First we recall the essence of the computation from [28]. Approaching the mirror dual of conifold singularity in the Kähler direction $t$, we know that the mirror map close to the singularity has the behavior
$\begin{split}t-t_{c}\sim u\log u\,.\end{split}$ (A.1)
Moreover, the Yukawa coupling $\kappa_{ttt}$ diverges like
$\begin{split}\kappa_{ttt}\sim{1\over u\log^{3}u}\sim{1\over(t-t_{c})\log^{2}(t-t_{c})}\,.\end{split}$ (A.2)
The question is how the GV invariants $\alpha^{n}_{0}$ must scale with $n$ to reproduce this behavior from the relation
$\begin{split}\kappa_{ttt}\sim\sum_{n=1}^{\infty}\alpha^{n}_{0}\,n^{3}\,e^{2\pi int}\,.\end{split}$ (A.3)
Making an ansatz
$\begin{split}\alpha^{n}_{0}\sim n^{\rho-3}\,(\log n)^{\sigma}\,e^{2\pi n{\rm Im}(t_{c})}\end{split}$ (A.4)
and approximating the sum over $n$ by an integral, one finds
$\begin{split}\kappa_{ttt}&\sim\sum_{n}n^{\rho}(\log n)^{\sigma}\,e^{-2\pi n{\rm Im}(t-t_{c})}\sim{\Gamma(1+\rho)\over(\Delta t)^{1+\rho}}(-\log\Delta t)^{\sigma}\end{split}$ (A.5)
with $\Delta t=2\pi{\rm Im}(t-t_{c})$. Comparing this to (A.2) one concludes that the ansatz works for $\rho=0$ and $\sigma=-2$. Hence, the GV invariants scale essentially exponentially $\alpha^{n}_{0}\sim\exp(2\pi n{\rm Im}(t_{c}))$ for large $n$ and we have verified that the divergence of the Yukawa couplings at the conifold singularity contains the information about this behavior.
This argument was for a single Kähler modulus, but should be generalized to more Kähler moduli. For the CY $\mathbb{P}_{1,1,1,6,9}[18]$ we have seen that there are two intersecting conifold loci so that we expect that the GV invariants $\alpha_{0}^{(n_{1},n_{2})}$ counting BPS 2-cycles with volume $n_{1}t_{1}+n_{2}t_{2}$ scale exponentially along both directions. Employing the software package CYTools [50, 51], we were determining the GV invariants up to total order $n_{1}+n_{2}\leq 200$ with the lowest ones listed in table 6.
Figure 6: Gopakumar-Vafa invariants of $\mathbb{P}_{1,1,1,6,9}[18]$.
However, these data were not yet sufficient to fix the precise form of the argument of the exponential function. Hence, also the potential appearance of the values (4.32) of the Kähler moduli $(t_{1},t_{2})$ at the point $P_{+}$ were not yet apparent.
For the other CY of interest, namely $\mathbb{P}_{1,1,2,2,6}[12]$ there is an intersection of a conifold locus and another degeneration locus at $t_{2}=0$. In table 7 we list the lowest GV invariants $\alpha_{0}^{(n_{1},n_{2})}$, which vanish for $n_{2}>n_{1}$ and satisfy the reflection property $\alpha_{0}^{(n_{1},n_{2})}=\alpha_{0}^{(n_{1},n_{1}-n_{2})}$.
Figure 7: Gopakumar-Vafa invariants of $\mathbb{P}_{1,1,2,2,6}[12]$.
In fact, we have computed all GV invariants up to total order $n_{1}+n_{2}\leq 200$. Here the situation turns out to be a bit clearer than for the previous CY. Inspection reveals that for fixed $n_{1}$ the GV invariants seem to follow a Gaussian distribution with a maximum at $n_{2}=[n_{1}/2]$ and width $\sigma^{2}\sim n_{1}$. Looking more closer, after making an educated ansatz, we determine the form of the GV invariants as
$\begin{split}\alpha_{0}^{(n_{1},n_{2})}\sim{e^{2\pi n_{1}\mu}\,{1\over n_{1}^{\rho+3}}\,\log^{\sigma}(n_{1})}\exp\left(-2\pi\lambda{(n_{2}-n_{1}/2)^{2}\over n_{1}}\right)\,\end{split}$ (A.6)
with $\mu=\lambda=4/3$ and $\sigma=0$. We notice that $\mu=4/3$ precisely matches the value (4.66) of ${\rm Im}(t_{1})$ at the degeneration point $P_{1}$. Moreover, for $n_{2}=0$ one can fit the GV data very well with $\rho=0$. Just sticking to the first column in 7, the exponential rate is $\mu-\lambda/4=1$, which happens to be the value (4.54) of ${\rm Im}(t_{1})$ at the emergent string limit $D_{\infty}\cap D_{1}$. Performing a similar computation as for the Quintic just for this column we would get the singular behavior
$\begin{split}Y_{t_{1}t_{1}t_{1}}\sim{1\over t_{1}-i}+\ldots\,,\end{split}$ (A.7)
which indeed matches (4.56) for the emergent string limit.
Taking also the other columns into account, the data reveal that $\rho$ is not constant but follows a plateau like behavior that in the following we approximate via the function
$\begin{split}\rho\approx{3\over 4}\tanh\left(\kappa{n_{2}(n_{1}-n_{2})\over n_{1}^{2}}\right)\,,\end{split}$ (A.8)
where $\kappa$ is a not too small parameter to guarantee a broad plateau. It is clear that the function can only depend on $n_{2}/n_{1}$ and must be invariant under the reflection $n_{2}\to n_{1}-n_{2}$. Following the previous computation for the quintic, one could try to estimate the four Yukawa couplings via the double sum
$\begin{split}Y_{t_{1}^{l}t_{2}^{3-l}}\sim\sum_{n_{1},n_{2}=1}^{\infty}n_{1}^{l}\,n_{2}^{3-l}\,\alpha_{0}^{(n_{1},n_{2})}\,e^{-2\pi(t_{1}n_{1}+t_{2}n_{2})}\end{split}$ (A.9)
with $l=0,\ldots,3$. Again, we can approximate the infinite sums via integrals over continuous variables $(n_{1},n_{2})\to(y_{1},y_{2})$. We did not succeed in analytically solving the appearing double integral. However, after introducing polar coordinates $y_{1}=r\sin\phi$, $y_{2}=r\cos\phi$ the integral over $r$ can be carried out analytically. The result can then be numerically integrated over the angle $40\leq\phi\leq\pi/$ 4 for various choices of $t_{1}$ and $t_{2}$ approaching zero. First, one realizes that (A.9) seems to diverge for $\Delta t=t_{1}-4/3+t_{2}/2$ to zero. The precise functional dependence cannot be uniquely fixed but we observe that the numerics is well consistent with a divergence
$\begin{split}Y_{t_{1}^{l}t_{2}^{3-l}}\sim{1\over\Delta t\log^{2}(\Delta t)}+\ldots\,.\end{split}$ (A.10)
This can be inferred from figure 8, which should then be a straight line.
Figure 8: Plots of $(\log^{2}(\Delta t)\,Y_{111})^{-M}$ as a function of $\Delta t$ for $\kappa=10$. Left $M=2/3$, middle $M=1$, right $M=3/2$.
From (4.65) we find at leading order
$\begin{split}\Delta t=t_{1}-t_{1}^{(0)}+t_{2}/2\sim x_{1}\log(x_{1})+\ldots\end{split}$ (A.11)
so that we realize that (A.10) is consistent with the leading divergence (4.74) found at the degeneration point $P_{1}$. In the same manner it is consistent with (4.85) for the point $P_{2}$ Having just this single example we cannot decide whether this is really the correct procedure but it made it evident that the asymptotic behavior of the GV invariants encodes information about the degeneration loci, which is the underlying structure of our proposed regularization procedure for the zero-point Yukawa couplings.
Appendix B Relation to singularities in Gauged Linear Sigma Models
In the proposed regularization method, it remains to better describe towards which singular point $t_{j,0}$ ($j=1,\ldots,h_{11}$) in Kähler moduli space the limit is taken. We observe that for all examples studied in this paper the CY manifolds admit a description in terms of Gauged Linear Sigma Models (GLSMs). The moduli space of these is parametrized by the complexified Fayet-Iliopoulos (FI) terms $\xi_{j}=\theta_{j}+ir_{j}$ for the $h_{11}$ $U(1)$ gauge fields. The classical moduli space encounters a singularity at the origin $\xi_{j}=0$, where all the different phases meet. However, due to quantum effects the FI-terms receive one-loop corrections, which change the location of the singularity. In fact, one can use the algorithm of [52] to determine the position of the quantum singularity.
Define $z_{j}=\exp(2\pi i\xi_{j})$, the quantum singular loci in the GLSM are at the solutions to the equations [52]
$\prod_{i=1}^{M}\big{\langle}\delta_{i}\big{\rangle}^{Q_{i,j}}=z_{j}\,,\qquad j=1,\ldots,h_{11}\;$ (B.1)
and
$\phantom{aa}\prod_{i\in I}\Big{\langle}\delta_{i}|_{H}\Big{\rangle}^{Q_{i,j}}=z_{j}\,,\qquad j=1,\ldots,{\rm rank}(H)\;,$ (B.2)
where $M$ denotes the number of $N=(2,2)$ chiral superfields and $\langle\delta_{i}\rangle$ the VEV of the operator $\delta_{i}=\sum_{j=1}^{h_{11}}Q_{i,j}\,\sigma_{j}$. Here $\sigma_{j}$ is the scalar part of the vector multiplet and $Q_{i,j}$ the $G=U(1)^{h_{11}}$ charges of the chiral superfields. Other components of singular loci are (B.2) and are obtained from subgroups $H\subset G$ such that the charges of the complementary set generate (with positive coefficients) all of $\mathbb{R}^{h_{11}-k}$ where $k$ is the rank of $H$. The product in (B.2) is over all chiral fields charged under $H$ and the $\delta_{i}|_{H}$ are obtained from $\delta_{i}$ by setting all scalar fields $\sigma$ related to the complement of $H$ to zero.
The procedure is the same for all models, so we will demonstrate it in the case of the two parameter model $\mathbb{P}_{1,1,1,6,9}[18]$ with gauge group $G=U(1)\times U(1)$. The gauge charges are
$Q=\begin{pmatrix}0&0&0&3&2&1&-6\\ 1&1&1&0&0&-3&0\end{pmatrix}\;.$ (B.3)
Inserting this into (B.1) and expressing the $\delta_{i}$ in terms of $\sigma_{j}$ gives
$z_{1}$ $=(-6s_{1})^{-6}(3s_{1})^{3}(2s_{1})^{2}(s_{1}-3s_{2})$ (B.4)
$z_{2}$ $=s_{2}^{3}(s_{1}-3s_{2})^{-3}\,,$
where for ease of notation $\langle\sigma_{j}\rangle=s_{j}$. These equations have solutions if
$6^{12}z_{1}^{3}z_{2}-(1-3^{3}2^{4}z_{1})^{3}=0\;.$ (B.5)
Moreover, the only subgroup of $G$ whose charges span $\mathbb{R}$ is the first $U(1)$ so that $H$ is equal to the second $U(1)$. Therefore (B.2) produces the additional relation
$z_{2}=s_{2}^{3}(-3s_{2})^{-3}=-1/27\;.$ (B.6)
Hence, the two relations (B.5) and (B.6) specify the quantum shift of the classical singularity at $\xi_{1}=\xi_{2}=0$. Comparing this with the singularity in the mirror dual complex structure moduli (4.22), we can read-off the identification
$\begin{split}z_{1}={1\over 3^{3}\cdot 2^{4}}\,\overline{x}\,,\qquad\quad z_{2}={1\over 3^{3}}\,\overline{y}\,.\end{split}$ (B.7)
This clearly shows that the intersection of the two conifold singularities corresponds to the quantum shifted singular point of the GLSM. Note that in the regime of large Kähler moduli, one has an expansion [53]
$\begin{split}q_{1}=z_{1}(1+O(z_{1},z_{2}))\,,\qquad\qquad q_{2}=z_{2}(1+O(z_{1},z_{2}))\,.\end{split}$ (B.8)
The same analysis also works for the Quintic and $\mathbb{P}_{1,1,2,2,6}[12]$.
Therefore, for the case that we have a description of the CY in terms of a GLSM we have learned that in a certain sense we are indeed taking a limit towards zero size, namely towards vanishing FI parameters $\xi_{i}=0$. Since due to quantum effects this is not really possible, the most natural thing to do is to take the limit towards their quantum corrected values.